### v1

# DiffHO: Diffusion-Based Synthesis of Hand-Object Interactions from Textual Descriptions

Abstract.
Generating natural hand-object interactions in 3D is challenging as the resulting hand and object motions are expected to be physically plausible and semantically meaningful. Furthermore, generalization to unseen objects is hindered by the limited scale of available hand-object interaction datasets. We propose DiffHO, a novel method to synthesize realistic, one or two-handed object interactions from provided text prompts and geometry of the object. The method introduces three techniques that enable effective learning from limited data. First, we decompose the task into a grasping stage and a text-based interaction stage and use separate diffusion models for each. In the grasping stage, the model only generates hand motions, whereas in the interaction phase both hand and object poses are synthesized. Second, we propose a compact representation that tightly couples hand and object poses. Third, we propose two different guidance schemes to allow more control of the generated motions: grasp guidance and detailed textual guidance. Grasp guidance takes a single target grasping pose and guides the diffusion model to reach this grasp at the end of the grasping stage, which provides control over the grasping pose. Given a grasping motion from this stage, multiple different actions can be prompted in the interaction phase. For textual guidance, we contribute comprehensive text descriptions to the GRAB dataset and show that they enable our method to have more fine-grained control over hand-object interactions. Our quantitative and qualitative evaluation demonstrates that the proposed method outperforms baseline methods and leads to natural hand-object motions. Moreover, we demonstrate the practicality of our framework by utilizing a hand pose estimate from an off-the-shelf pose estimator for guidance, and then sampling multiple different actions in the interaction stage. The project page along with videos can be found at https://diffh2o.github.io/
1. Introduction
Modeling and understanding hand-object interactions is an important and crucial task to empower machines to interact with and assist humans. Being able to seamlessly generate hand-object motions holds the promise to enable synthetic data generation (Leonardi et al., 2023), assist robots in training human-robot interactions in simulation, or enhance realism in virtual manipulation experiences. For example, in virtual reality (VR), hand interactions still often rely on heuristics and controllers that simply attach objects to the hand according to pre-defined grasps. Being able to faithfully produce object manipulation based on an input signal such as text or a few past frames of hand and object poses could largely increase the immersiveness of such interactions.
Generating realistic hand-object interactions in 3D is challenging as the resulting motions are required to be plausible in several different aspects. First, the motions must be plausible in terms of geometry, such hand and object intersections are minimized and the grasp appears stable. Second, the motions must be plausible in terms of semantics, such as the hands respect natural object affordances (e.g., a cup is grasped by its handle and not flipped upside down). Third, the motions must be plausible in terms of time, such as the hand and object motions are synchronized and the dynamics appear natural. Another challenge in generating hand-object interactions comes from the limited scale of existing hand-object datasets, which are around 10x smaller than human motion datasets (Mahmood et al., 2019) and 1000x smaller than image datasets (Deng et al., 2009).
Several recent works have successfully leveraged diffusion models for full-body human motion generation but either neglect objects (Tevet et al., 2023; Karunratanakul et al., 2023) or focus only on coarse motion of larger objects (e.g., a chair) without finger predictions (Xu et al., 2023). Another method, IMoS (Ghosh et al., 2023), can generate hand-object interactions with smaller objects (e.g., a cup or stapler) by first optimizing for upper body motions and then optimizing for object poses in a post-processing step. However, IMoS assumes that the object is in hand from the beginning, can handle only objects seen during training, and, as shown in our qualitative evaluation, the optimization often yields artifacts (e.g., hand-object interpenetration or non-plausible contact points). In this paper, we aim to learn a model that can generate natural, fine-grained hand-object interactions and generalize to objects unseen during training.
Towards this goal, we contribute a novel diffusion-based framework, DiffHO, that can generate hand-object interactions based on provided text prompts and geometry of the object. We propose several techniques to deal with the challenges of data scarcity and object generalization. Decoupling Grasping and Interaction: Hand-object interactions can be invariably split into a grasping phase where the hand(s) approaches the static object for grasping, and an interaction phase where the object is manipulated based on an action intent. For example, drinking from a cup can be performed with multiple grasps, and given a grasp, multiple actions are possible. We use this simple observation to obtain a grasping and an interaction diffusion model, and introduce an inpainting technique called subsequence guidance to allow continuity between the two outputs. Different from previous diffusion frameworks that employ coarse-to-fine (2D trajectory-to-full motion) two-stage generation (Karunratanakul et al., 2023), we split the task temporally. Pose Representation: We introduce a compact representation that models hands in parametric space. To couple the hands to the object, we represent the global hand positions relative to the normalized object pose at the initial frame and further include distance information between the 3D hand joints and the object surface. Unlike previous representations utilized in human motion diffusion (Tevet et al., 2023) and hand-object interaction (Christen et al., 2022), we find a non-redundant representation directly in parametric space to perform best. Controllability: We provide two ways to add more control to our diffusion models’ output. First, we propose control through fine-grained textual descriptions. To this end, we contribute textual descriptions to the GRAB dataset and show that they enable more robustness to unseen test prompts and increase controllability, e.g., by allowing to define the interacting hand and the action to perform through the input prompts. Second, in grasp guidance, we leverage a single grasp reference which connects the grasping and interaction phase and use it to guide our two diffusion models at inference time. This grasp pose can either be obtained from motion capture or from an image-based pose estimator (Pavlakos et al., 2023) as we show in our experiments Section 5.8.
To demonstrate the effectiveness of our approach, we run perceptual studies and experiments on the GRAB dataset (Taheri et al., 2020). We first compare against IMoS (Ghosh et al., 2023). Our results indicate that our method generates better results across a variety of physics-based and motion metrics. To justify our technical contributions, we compare our final model against human-body diffusion baselines by adapting these methods to hand-object interactions. We show that we generate motions of higher quality while enabling more control of the outputs. Furthermore, we demonstrate that detailed textual descriptions increase robustness to unseen texts and offer better control of the diffusion model than training with heuristics-generated text descriptions. Decoupling and grasp-guidance brings high practical value to DiffHO. We show this with the first application of using an image-based pose/shape estimator (Pavlakos et al., 2023) to guide the diffusion to the grasp, and thereafter sampling multiple different actions through text. We prove the utility of our method by largely outperforming the state-of-the-art on a standard benchmark and in perceptual studies.
Our contributions can be summarized as follows:
To the best of our knowledge, DiffH2O is the first method that synthesizes hand-object interactions on unseen objects from textual descriptions, unlike previous generative models for human-body motion and seen-object HOI synthesis.
We propose a two-stage diffusion process that splits generation into grasping and action-based interaction.
To increase the controllability of outputs and improve generalization to unseen objects, we propose grasp guidance and subsequence guidance that can be applied at inference time to the diffusion process.
We contribute detailed textual annotations to the GRAB dataset along with experiments showing that these textual descriptions increase robustness and enable fine-grained control of the model output.
Code and data will be made public upon acceptance.
2. Related Work
2.1. Hand-Object Synthesis
Recent efforts in interaction synthesis have largely been driven by the surge of high-quality human-object interaction datasets (Taheri et al., 2020; Brahmbhatt et al., 2019; Hampali et al., 2020; Brahmbhatt et al., 2020; Chao et al., 2021; Kwon et al., 2021; Fan et al., 2023; Liu et al., 2022). Some studies focus on generating coarse full-body object interactions (Zhang et al., 2022a; Wang et al., 2021; Lee and Joo, 2023; Hassan et al., 2023; Luo et al., 2021), such as carrying or moving around boxes. FLEX (Tendulkar et al., 2023) trains a hand and body pose prior and later optimizes the priors to achieve diverse, static full-body grasps. GOAL (Taheri et al., 2022) and SAGA (Wu et al., 2022) use CVAEs to generate approaching motions for full-body grasps, whereas TOHO (Li et al., 2024) models both approaching and manipulation tasks via neural implicit representations. Similarly, Braun et al. (2024) model the full range of motion, leveraging physics simulation and reinforcement learning. In contrast to these works, we focus on fine-grained hand-object interactions and generalization to unseen objects. Furthermore, we leverage language prompts as input to generate interaction sequences. Among methods that model full-body object interaction, closest to our work is IMoS (Ghosh et al., 2023), a two-stage method to generate hand-object interactions on seen objects based on language commands. Starting from a grasping state, they first generate body motions and then optimize for object trajectories using a heuristics-based optimization to model hand-object interactions. By contrast, our model directly predicts the hand and object poses, models both approaching and manipulation phases, and generalizes to unseen objects.
Another line of research focuses on generation of hand-object interaction sequences in isolation from full-body due to wide applications in VR and the need for a dedicated model for fine-grained details of finger motion. One prominent solution is to turn to physical simulation and reinforcement learning (Christen et al., 2022; Mandikal and Grauman, 2021; Garcia-Hernando et al., 2020; Rajeswaran et al., 2018; Qin et al., 2021). A number of studies propose to learn dexterous manipulation tasks from full human demonstration data either collected via teleoperation (Rajeswaran et al., 2018) or from videos (Garcia-Hernando et al., 2020; Qin et al., 2021). Mandikal and Grauman (2021) propose a reward function that incentivizes dexterous robotic hand policies to grasp in the affordance region of objects. She et al. (2022) propose a new dynamic state representation to generate dexterous grasps. D-Grasp (Christen et al., 2022) leverages reinforcement learning and physics simulation to generate diverse hand-object interactions from sparse reference inputs. ArtiGrasp (Zhang et al., 2024) extends this to two-handed grasps and generates articulated object motions.
To model hand-object interaction sequences, another solution is to rely on purely data-driven frameworks. For example, (Zhou et al., 2022; Anonymous, 2023) propose methods that enable denoising hand-poses from noisy sequences of hand-object poses. Ye and Liu (2012) predict the local hand pose given full body and object motion. Similarly, ManipNet (Zhang et al., 2021) predicts local hand poses for two-handed interactions based on wrist-object trajectories. Given object motion of articulated objects, CAMS (Zheng et al., 2023) predicts one-handed poses that align with the object motion, whereas Chen et al. (2023b) focus on grasping objects using dexterity in the environment. All these works either assume a hand-object trajectory, focus only on one hand, ignore semantic action information, or do not generalize well to unseen objects. In contrast, we propose a framework that allows the synthesis of two-handed object interactions from text as well as generalization to unseen objects.
2.2. Diffusion for Motion Synthesis
Diffusion models (Sohl-Dickstein et al., 2015) have gained popularity in many domains, such as image generation (Ho et al., 2020), video generation (Yang et al., 2023), audio synthesis (Kong et al., 2021) or hand reconstruction (Ye et al., 2023b, a). More recently, diffusion models have also been adapted to human motion synthesis (Tevet et al., 2023; Zhang et al., 2022b). Various improvements have been proposed, such as integrating physics (Yuan et al., 2023), scene-awareness (Huang et al., 2023), increasing efficiency by diffusing in a pre-trained latent space (Chen et al., 2023a), or the composition of multiple actions (Athanasiou et al., 2022). To enable more controllability during motion synthesis, a two-stage framework, GMD, that guides the diffusion model towards target objectives at inference time is proposed in Karunratanakul et al. (2023). However, GMD contains design choices specific to human motion generation, by first generating a 2D root trajectory over the whole sequence, and then generating corresponding full-body poses, which is not directly transferable to HOI. While these works focus on human motion in isolation from objects, InterDiff (Xu et al., 2023) generates human-object interactions via diffusion models. Their work is improved upon by recent concurrent works (Diller and Dai, 2023; Li et al., 2023; Peng et al., 2023) that leverage contact-based predictions in combination with inference time-guidance to improve the interaction quality. However, these works focus on full-body motions and neglect intricate hand-object interactions. In contrast, we focus on the detailed interactions of fingers and objects. Most similar to ours, the concurrent work MACS (Shimada et al., 2023) proposes a diffusion model for hand-object motion synthesis. Their focus is on synthesizing interactions for a single object with varying mass, while our focus is and generalization to unseen shapes and interaction synthesis conditioned on text.
3. Denoising Diffusion Probabilistic Models
Denoising Diffusion Probabilistic Models (DDPM) (Sohl-Dickstein et al., 2015; Ho et al., 2020) aim to model the probability distribution of a given dataset and can model complex distributions of images (Rombach et al., 2022), videos (Ho et al., 2022) or time-series data such as motion sequences (Tevet et al., 2023).
The diffusion model involves a forward process, which is a Markov chain and consists of sequentially adding Gaussian noise to the data . Each step in the forward noising process results in a distribution, , where is a hyper-parameter and , where is the final number of steps. Alternatively, can be expressed in terms of as
where, .
The reverse process, also called the denoising process is another Markov chain with learned Gaussian transitions starting at . The denoised distribution at step is given by, , where are the model parameters to be learnt, is the mean, and is the variance. The process of data generation involves sequentially drawing samples from this distribution. The variance is usually set to and which is implemented using an auto-encoder trained to match the mean of forward process posterior which is another Gaussian represented as, , where
and .
The auto-encoder could directly estimate the mean or in Eq. 2, or the noise which is used in the noising process of Eq. 1. When estimating the noise, the denoised mean is obtained as , where is the auto-encoder output. The auto-encoder is trained to minimize
where, is either , , or .
Conditional diffusion models generate output conditioned on various kinds of input such as text (Rombach et al., 2022) and other semantic information (Zhang et al., 2023). Such a model is obtained by providing the condition as additional input to the auto-encoder resulting in outputs , or , where c is the conditional input encoding.
3.1. Guiding the diffusion model
Diffusion models can be guided to generate outputs of the required form. Two popular methods to guide the diffusion models are classifier guidance (Dhariwal and Nichol, 2021) and classifier-free guidance (Ho and Salimans, 2022). We describe here the classifier-guidance method which is more relevant to our work and refer the reader to (Ho and Salimans, 2022) for details on classifier-free guidance.
The classifier-guidance method approximates , where c is the guidance, by a Gaussian:
Intuitively, the scaled gradient of the guidance signal, is used to nudge the denoised mean to generate samples in accordance with the guidance c. could represent the class probability which can be implemented using a neural-net (Dhariwal and Nichol, 2021), or could represent a cost term that is optimized (Karunratanakul et al., 2023). The classifier guidance method has the advantage that any guidance signal that is differentiable can be used to guide the diffusion and can use pre-trained diffusion models without retraining them.
4. Method
We focus on the task of generating a sequence of hand and object poses , where indicate a sequence of left and right hand poses, and is a sequence of object poses given some input conditioning signal c, which in our case comprises a text description and the object mesh. In other words, we want to model a conditional probability distribution with a motion DDPM. The optional term is a differentiable goal function that should be minimized, e.g., the distance to a target grasp. See Fig. 2 for an overview.
In Sec. 4.1, we introduce a compact representation that tightly couples the hand and object poses and helps in generating realistic hand-object interactions. We propose a two-stage generation of hand-object interactions in Sec. 4.2 and grasp guidance in Sec. 4.3. Lastly, in Sec. 4.4, we introduce our newly-collected detailed textual annotations for the GRAB dataset that enable more controlled generation of hand-object interactions.
4.1. Canonicalized Hand-Object Representation
Our proposed representation of length N contains information about both hands and an object. In particular, the object poses are defined as , where is a sequence of 3D object locations, with , and is a sequence of 3D object orientations, with given in 6D representation (Zhou et al., 2019). The hands are represented by: , where denotes the left and denotes the right hand. This representation is based on the parametric MANO hand model (Romero et al., 2017). Specifically, the global 3D positions of the hands are given by and the global 3D orientations are represented by . To achieve more natural poses, we define the local hand pose in a PCA space of the MANO model with the first 24 components as . Lastly, denotes distances between each hand joint (including the fingertips) and its closest point on the object mesh. As we show in our ablations (see Tab. 4) this representation reduces physical artifacts such as interpenetration. We first normalize the object positions to be at the origin and then adjust the hands’ positions accordingly. Hence, all position components are relative to the object origin at the initial frame. Due to the limited data, such a representation performs better than a per-frame object-relative representation (Christen et al., 2022; Zhang et al., 2024) as we show in 5.6.
4.2. Two-Stage Hand-Object Generation
We propose to split the synthesis into grasping and interaction stages to facilitate generalization to unseen objects. The grasping stage models the motion from an initial pose to an object grasp. The interaction stage performs an action-based manipulation of the object. Our two-phase approach is motivated by the fact that any hand-object interaction involves an approaching phase where one or two hands grasp a static object, followed by an intent-driven interaction with the object. By decoupling the action-based manipulation from grasping, we can leverage the entire set of motions in the dataset for training the grasping phase, irrespective of the action that is performed in the interaction phase. This insight helps counteract the limited scale of the dataset.
As shown in Fig. 2, our grasping and interaction diffusion models are conditioned on three types of inputs to encode intent, namely, 1) text prompts describing the object and the action to be carried out, 2) object mesh which provides the geometry of the object, and 3) the time step in the denoising process. We use CLIP (Radford et al., 2021) embeddings to encode the text prompt denoted by , BPS (Prokudin et al., 2019) to encode the object mesh denoted by and an MLP to encode the time step as in previous works (Karunratanakul et al., 2023; Dhariwal and Nichol, 2021). We provide more implementation details in supplementary material and describe the grasping and interaction phases below in detail.
4.2.1. Grasping Stage
The grasping phase is defined as a sequence containing one or two hands approaching a static object from a rest pose until the object is grasped. As the object is static throughout this phase we do not predict the object motion and only model the hand pose sequence. In order to obtain the training data for this phase we use heuristics based on object poses to determine sequence boundaries from the larger action sequences and normalize the sequence lengths by interpolating or downsampling the poses. The text description in this phase is a generic sentence (”The person grasps the ¡object¿.”).
We use the diffusion model that directly predicts the noise (see Section 3) in classifier-free manner (Ho and Salimans, 2022) which is more susceptible to inference-time guidance as shown in GMD (Karunratanakul et al., 2023). In other words, our grasping diffusion model training loss is given by
where indicates the synthesized motion at diffusion step .
4.2.2. Interaction Stage
This phase comprises the motion of both hands and the object after a grasp has been established. In this phase, the hands manipulate and interact with the object according to the action defined through the textual prompt. We obtain the training data for this phase in a similar way as the grasping phase based on object pose heuristics. We employ a diffusion model in a classifier-free manner (see Section 3), i.e., we directly predict the denoised output. This is shown to result in less noisy, higher quality motion than (but less susceptibility to guidance) (Karunratanakul et al., 2023), which is critical during the interaction phase where the hand-object interaction takes place. Our training loss is given by
4.2.3. Grasp to Interaction Stage Transition
Maintaining a seamless transition between the grasping and interaction outputs is critical to generating realistic motion sequences. We rely on inference-time guidance (Karunratanakul et al., 2023) and motion imputing on the interaction model to achieve a smooth transition. To this end, the interaction diffusion model is trained to generate the complete sequence (including grasping). We then impute the entire sequence generated by the (guided) grasping model to the start of the interaction sequence and apply guidance. To perform the imputation we define respectively the projections and that resize the grasp reference or sub-sequence poses to that of by filling in zeros. Similarly, and represents the imputation regions of the grasp reference and sub-sequence on . At each denoising step, the imputated denoised output, , is given by
where and represents elementwise multiplication. We use the classifier guidance mechanism described in Sec. 3.1 and approximate the gradient signal for sub-sequence guidance as
where denotes the set of grasping poses and denotes the generated grasping poses from the grasping diffusion model. We show in Sec. 5.6 that this technique leads to a smoother transition between pre- and interaction compared to a single frame transition.
4.3. Grasp Guidance
To add controllability to the generation and improve generalization to unseen objects, we introduce grasp guidance for the grasping phase, which can be optionally be added at inference time. We leverage a single grasp frame, defined as the last frame of the grasping phase, that provides a prior about how and with which hand an object should roughly be grasped. Such a grasp can be obtained from motion capture or from an off-the-shelf hand-object pose estimator (Pavlakos et al., 2023) as demonstrated in Sec. 5.8. We provide inference-time guidance to our trained diffusion model to grasp the object in a pre-defined manner. More specifically, we approximate the gradient of the guidance signal in Eq. 4 as
where is the hand(s) pose in the last frame of the grasping sequence, and is the pre-defined grasp pose.
| Method | Backbone | SD [] () | OD []() | IV [] () | ID [] () | CR | MM () | DIV () |
Real Mocap (GRAB) | - | - | 0.183 | 5.65 | 5.89 | 0.08 | 0.22 | 1.06 | 0.708 | 0.729 | ||||
| IMoS (Ghosh et al., 2023) | CVAE | 0.002 | 0.149 | 7.14 | 11.47 | 0.05 | 0.25 | 1.02 | 0.579 | 0.588 | |||||
| DiffHO | Transformer | 0.088 | 0.185 | 6.65 | 8.39 | 0.067 | 0.33 | 1.09 | 0.760 | 0.810 | |||||
| DiffHO | UNet | 0.109 | 0.188 | 6.02 | 7.92 | 0.064 | 0.30 | 1.08 | 0.833 | 0.875 | |||||
Real Mocap (GRAB) | - | - | 0.183 | 3.40 | 5.63 | 0.071 | 0.20 | 1.08 | 0.683 | 0.784 | ||||
| IMoS* (Ghosh et al., 2023) | CVAE | 0.002 | 0.132 | 10.38 | 12.45 | 0.048 | 0.22 | 1.05 | 0.561 | 0.581 | |||||
| DiffHO | Transformer | 0.133 | 0.185 | 7.99 | 10.87 | 0.073 | 0.24 | 1.09 | 0.750 | 0.803 | |||||
| DiffHO | UNet | 0.134 | 0.179 | 9.03 | 11.39 | 0.086 | 0.23 | 1.09 | 0.755 | 0.837 |
4.4. Textual Augmentations
Our model generates hand-object interactions based on textual inputs. To the best of our knowledge, there is currently no hand-object interaction dataset that contains detailed textual descriptions. For instance, the GRAB dataset (Taheri et al., 2020) only provides categorical action labels. To address this limitation, we auto-generated sentences using the template “the person ¡verb¿ + ¡object¿” (Ghosh et al., 2023) (referred to as “simple” text). However, these sentences lack detailed information, which limits the controllability of the model via textual inputs. Therefore, we contribute carefully annotated textual descriptions of the GRAB dataset (referred to as “detailed” sentences). We instructed our annotators to carefully watch each ground truth video and describe the actions in three distinct stages of pre-action, action, and post-action. These descriptions are further augmented with hand and positional information. An example from the dataset is as follows: ”The person picks up the apple with the right hand, passes it to their left hand, and then places it back with their right hand”. These detailed descriptions enhance the dataset, enabling more accurate evaluation and precise control of hand-object interactions as we show in Sec. 5.7.
5. Experiments and Results
5.1. Baselines
Our method focuses on synthesizing detailed hand-object motion based on text input. To the best of our knowledge, DiffHO is the first method to tackle this problem. While there is no direct baseline for hand-object motion synthesis based on language, the closest to our approach is IMoS (Ghosh et al., 2023), which focuses on text-based whole-body human-object interaction synthesis. IMoS first generates a body articulation based on an instruction label (action + object), and then optimizes for the object pose using a grasp heuristic. Since our paper focuses on hand articulations and object motions, we omit the full-body movements from our evaluations.
To compare our method against human-body diffusion models, we adopted these methods to the HOI setting. All variants use our proposed representation. MDM (Tevet et al., 2023) does not use guidance, and for GMD (Karunratanakul et al., 2023) we use gradients to guide the model towards the grasp frame, but omit the 2D-trajectory generation stage because it is specific to human motions.
5.2. Data
GRAB: We utilize the subject-based split of the GRAB dataset (Taheri et al., 2020) proposed in IMoS to run a direct comparison. However, as this split does not contain unseen objects, we also create a new split featuring unseen objects based on the criteria of object similarity and semantics (see supplemental material for details).
HO-3D: To further evaluate guidance and object generalization, we estimate hand-object poses on the HO-3D dataset (Hampali et al., 2020) and guide our model to an estimated target pose.
5.3. Evaluation Metrics
Following previous works on motion synthesis (Ghosh et al., 2023; Tendulkar et al., 2023; Taheri et al., 2022; Braun et al., 2024), we report several physics-based metrics and evaluate our model’s accuracy through metrics that assess motion diversity and action features.
Physics Based Metrics: We report the interpenetration volume (IV) as the amount of MANO vertices that penetrate the object mesh and the maximum interpenetration depth (ID). Furthermore, we compute the contact ratio (CR) as the average ratio of hand vertices that are in contact, which is defined as points within a 5mm threshold of the object mesh.
Motion Diversity Metrics: We report the sample diversity (SD) between wrist trajectories in motion space. We sample the same input conditions five times and compute the pairwise Euclidean distance between the samples. We then report the mean sample diversity over all test prompts. We also provide the overall diversity (OD), which measures the pairwise Euclidean distance between all test samples. Since the wrist positions are normalized with respect to the initial object pose, no further normalization is applied.
Guidance Metrics: To assess how well the models adhere to the guidance signal, we report the mean error between the predictions and the reference grasp (GE), the ratio at which the diffusion model maintains the correct handedness according to the reference (HA). Furthermore, we evaluate the wrist velocities at the transition between grasping and interaction phase ().
Action Feature Based Metrics: To assess the naturalness and variety of the generated motions, we use the action recognition accuracy, diversity, and multimodality metrics, following (Ghosh et al., 2023). We provide additional metrics (e.g. FID, KID), standard deviations and details about how these metrics are computed in our supplemental material.
5.4. Comparison to the State-of-the-Art
We conduct a set of experiments to compare to relevant baselines quantitatively. To demonstrate the generality of our method, we present qualitative examples of our method in Fig. 3 and supplementary materials (along with failure cases).
Setting 1 - Interaction Only: We compare our approach against the state-of-the-art human-object motion synthesis approach of IMoS in their postgrasp-only setting and report results in Tab. 1. To ensure a fair comparison in this experiment, no grasping and neither grasp guidance nor our detailed text annotations are used. We test our hand-object representation with two backbones, MDM’s transformer (Tevet et al., 2023) and GMD’s UNet (Karunratanakul et al., 2023). Our model significantly outperforms IMoS across all physics and motion diversity metrics as well as on the action recognition accuracies. In particular, our method yields motions that are significantly more diverse (higher SD and OD), exhibit less interpenetration (lower IV and ID), and encode motion features that align with action type (higher action recognition accuracies). Note that IMoS assumes the availability of a ground-truth test motion to initialize the first few frames with it, while ours predicts the entire sequence. Yet, our approach achieves considerable performance improvements. While IMoS provides compelling realistic full body synthesis results, fine-grained finger motions as well as post-optimized object motions lag behind the realism and diversity demonstrated by our method, as shown in Fig. 3 and in our supplementary video. The higher scores of DiffHO compared to real mocap from the GRAB dataset in Tab. 1 can occur because the action classifier is trained on GRAB’s training-split, which is the distribution our model covers,whereas some of the test sequences in GRAB (real mocap) have quite distinct motions that the classifier may misclassify. However, the high scores demonstrate the representative power of DiffHO for generating realistic hand-object sequences that capture inherent motion and action semantics. Our method with a UNet backbone outperforms a transformer backbone in most metrics, which is in line with GMD’s findings that UNet works better in low-data regimes. To assess the action recognition accuracy, we train two models, one that uses only hand 3D joints as input and another incorporating object 3D position as additional input. We demonstrate that DiffHO’s synthesized object motions contribute to larger performance improvements compared to the post-optimized motions computed by IMoS ( improvement for ours vs for IMoS) as shown in Tab. 1. This confirms the benefits of jointly modeling hand-object motion instead of decoupling it into separate components.
| Method | IV [] () | GE | HA | |
| MDM | 8.98 | 0.38 | 0.45 | 0.14 |
| GMD | 9.03 | 0.49 | 0.46 | 12.68 |
| Ours | 7.40 | 0.12 | 0.87 | 0.23 |
Setting 2 - Grasping + Interaction: In this experiment, we switch to the setting of generating the whole sequence of grasping and interaction, and are interested how well the models can be guided while generating high quality motions (without physical artifacts). We report results in Tab. 2 and find that our method outperforms baselines from human motion diffusion. In particular, our method better adheres to the grasp frame in terms of grasp error (GE) and handedness accuracy (HA), and achieves more natural grasps (lower IV). This justifies our contributions of two-stage separation and subsequence guidance over previous diffusion models.
5.5. Perceptual Study
To evaluate the visual quality of our motions, we conducted a perceptual user study on a set of 21 participants, who were randomly assigned one of four sets, each consisting of 30 randomly selected sequences. We compared our approach against IMoS. We displayed results from our method and IMoS side-by-side in random order, along with the input sentence description used to synthesize motions. We asked the following questions to the participants: “Which sequence is more realistic?” and “Which sequence has more more variety in motion?”. We defined a sequence to be more realistic if the overall motion looks human-like, e.g., if the objects are grasped realistically and there are less artifacts such as floating objects or interpenetration. Moreover, a sequence has more variety if the object is manipulated multiple times and in distinct ways. In 63.1 of the responses, participants selected our method as the most realistic compared to IMoS. Even more distinctively, in 72.9 of the responses, participants favored our method to contain more variety than IMoS. We obtained statistically significant Fleiss’ kappa scores (Fleiss, 1971) of 0.34 and 0.43 for Realism and Diversity, respectively, at p¡0.05. Please see supplemental material for more details.
5.6. Ablation Studies
We perform several ablations to justify our technical contributions. To evaluate the controllability gained through guidance, we compare our final method with several variants in Tab. 3. We use a base model which indicates a single stage model without guidance. We then gradually add our contributions, first the grasping and interaction task separation (2-ST), the grasp guidance (GG) and the sub-sequence guidance in the interaction phase (SS) as described in Sec. 4.2. As expected, the grasp reference is being ignored without guidance (top 2 rows in Tab. 3), leading to high grasp errors and low hand correctness. Adding guidance allows for better performance in grasp error and hand accuracy, however, a high indicates a sudden jump of the hands between grasping and interaction. To mitigate this, we introduce the smooth transition technique that leads to the best overall score at the cost of a slight increase in grasp error and decrease in hand correctness. The interpenetration volume is lowest with our final model, indicating the best generalization to unseen objects. To validate our canonicalized representation, we compare against a variant inspired by D-Grasp (Christen et al., 2022), our representation without SDF, using object-relative positions at every step, and finally our representation of object-relative with regards to the initial frame. Our final representation achieves less interpenetration and better contact modeling.
| Component | Metrics | |||||
| 2-ST | KF | SS | IV [] () | KE | HA | |
| 7.48 | 0.31 | 0.52 | 0.21 | |||
| ✓ | 9.03 | 0.49 | 0.46 | 12.68 | ||
| ✓ | ✓ | 8.52 | 0.06 | 0.97 | 5.37 | |
| ✓ | ✓ | ✓ | 7.40 | 0.12 | 0.87 | 0.23 |
| Pose Representation | Metrics | ||
| IV [] () | ID | CR | |
| D-Grasp (Christen et al., 2022) | 5.56 | 8.60 | 0.055 |
| Ours w/o SDF | 10.79 | 11.42 | 0.073 |
| Ours w/ object-relative pose | 9.38 | 11.41 | 0.077 |
| Ours w/ first frame relative pose (Ours) | 9.03 | 11.39 | 0.086 |
5.7. Controllability via Text Prompts
In this experiment, we train DiffHO once with our detailed text labels and once with simple text prompts generated from the GRAB objects and action labels. As shown in Tab. 5, our textual descriptions offer significant advantage by enabling control over handedness. To gauge controllability, we measure whether the active hand in the generated motion corresponds to prompted hand. Our model, trained with detailed text annotations achieves a correct overall handedness accuracy of 86.5% against 59.3% accuracy when trained with simple text prompts. We see a performance drop when training on detailed texts and testing on simple texts (0.731) and vice-versa, which was similarly observed in (Shvetsova et al., 2023) due to a style gap. Similarly, we measure action correctness by evaluating the match ratio between the prompted action and the action demonstrated in the generated motion. Using detailed textual descriptions enables our model to generate motions that align more accurately with the text input, demonstrated by a 93.5% accuracy compared to the 54.8% accuracy achieved by our model tested with simple text descriptions. Since the test inputs are much more diverse with the detailed textual descriptions, this furthermore demonstrates the robustness to unseen sentences. We also report the cosine similarity between the different sets of texts as a reference.
| Train | Test | Action | Hand correctness | Cosine | |||||
| Input | Input | Correctness | Right | Left | Both | Total | Similarity | ||
0.897 | n/a | n/a | n/a | n/a | 0.72 | ||
0.731 | n/a | n/a | n/a | n/a | 0.43 | ||
0.548 | 0.709 | 0.111 | 0.0 | 0.593 | 0.43 | ||
0.935 | 0.869 | 0.862 | 0.75 | 0.865 | 0.66 | ||
5.8. Motion Generation from Estimated Pose
We run an image-based pose estimator (Pavlakos et al., 2023) on single images from HO3D (Hampali et al., 2020) to obtain object-relative hand poses. We then use these hand poses as reference grasps and pass them through our diffusion framework together with multiple different textual descriptions. This highlights the practicality of our framework-given a single grasp reference, we can generate multiple diverse sequences.
6. Discussion and Conclusion
In this work, we have introduced a framework to generate plausible hand-object interactions from textual descriptions. Specifically, we have proposed a two-stage diffusion method that separates sequences into a grasping and interaction stage and uses single keyframe grasps as guidance to the diffusion model to improve generalization across unseen objects. Through our compact representation, hands are coupled to object motions resulting in realistic interactions. While we provide a first step towards object generalization, there are limitations of our framework. To remove physical artifacts, physics could be explicitly integrated into the diffusion process (Yuan et al., 2023). The inference time is still relatively slow in diffusion models. To increase efficiency, operating in latent space is an interesting direction to explore (Chen et al., 2023a). Lastly, grasp references for guidance which are far away from the training distribution tend to be ignored by our model. Therefore, exploring new ways to guide the diffusion model towards out-of-distribution samples is a crucial future research direction.
Appendix - Supplementary Material
Appendix A Implementation Details
DiffHO architecture is based on UNet with Adaptive Group Normalization (AdaGN) which was originally proposed in (Dhariwal and Nichol, 2021) and also adapted in (Karunratanakul et al., 2023) for sequence prediction tasks. We adapt this network architecture for our two-stage design, that involves pre-grasp and post-grasp diffusion models. We provide the hyperparameter settings used in our architecture in Table 6.
Appendix B Experimental Details
B.1. Training the action classifier.
We train a standard RNN action recognition classifier on the GRAB dataset and use the final layer of the classifier as the motion feature extractor for calculating action recognition accuracies as well as diversity (DIV) and multimodality (MM) scores, following (Ghosh et al., 2023) and (Guo et al., 2020). We employ the online available code from IMoS and use the same network architecture and parameters to train our model.
B.2. Unseen object split
To assess our model’s ability to generalize to unseen classes, we also introduce a new split comprising unseen objects that were not present during training. Our test set is composed of the following objects: “apple, mug, train, elephant, alarm clock, small pyramid, medium cylinder and large torus”. Note that the training set includes pyramids, cylinders, and torus in other sizes. We exclude small pyramids, medium cylinders, and large torus from the training set to test our models’ ability to generalize to different sizes of the same objects
B.3. Details of the Action Recognition Evaluation Metrics.
Recognition Accuracy.
Given a pretrained classifier, we input motions generated by DiffHO as well as our baselines and report top-1 percentage accuracy. The recognition accuracy indicates the correlation of the action type and the motion.
Diversity.
We compute the diversity score by computing the variance of the features extracted from the action classifier across all action categories. Given a set of features extracted from generated motions across all action categories, two subsets with the same size, N, are randomly sampled. These features are denoted as for the first subset and for the second subset. With , the diversity is defined as
Multimodality.
Different from the diversity metric, multimodality computes variance only within a specific action. An overall score is attained based on averaging the variances across all action types.
Fréchet Inception Distance (FID).
FID is computed based on the (Fréchet) distance of the features of real ground-truth motions and generated motions. It has the assumption that the features computed from the real data and synthesized data have a Gaussian distribution. Earlier work demonstrated that for small sample sizes, FID metric is biased and not stable.
Kernel Inception Distance (KID).
KID metric relaxes the Gaussianity assumption and aims to improve upon FID. It measures the squared Maximum Mean Discrepancy (MMD) between the feature representations of the real and generated samples using a polynomial kernel. This metric has been found to be more robust when the sample size is small (Noguchi and Harada, 2019).
| Pre-Grasp | Post-grasp | |
| Parameter | Model | Model |
| Batch size | 32 | 32 |
| Base channels | 256 | 512 |
| Latent dimension | 256 | 512 |
| Channel multipliers | (1, 1, 1) | (2, 2, 2, 2) |
| scheduler | Cosine (Nichol and Dhariwal, 2021) | |
| Learning rate | 1e-4 | 1e-4 |
| Optimizer | AdamW (wd = 1e-2) | |
| Training | 400 | |
| Diffusion loss | prediction | prediction |
| Diffusion var. | Fixed small | |
| Model avg. beta | 0.9999 | |
B.4. Visualization Pipeline.
We use AITViewer (Kaufmann et al., 2022) to visualize motions synthesized by DiffHOas well as our baselines. IMoS synthesizes 15 frames of human-object interactions. They determine the hand which is in contact with the object according to GT, which can lead to a jump of the object from one hand to the other for certain actions. To make the frame rates comparable with DiffHO, we upsample the synthesis results of IMoS to 75 frames using spherical interpolation. Therefore, this jump may result in an artifact in which the object is floating from one hand to another.
Appendix C Additional Experiments
C.1. Action Recognition Experiments
Training action classifier on the whole GRAB data.
IMoS (Ghosh et al., 2023) reports results using the whole dataset to train the action classifier on ground-truth motions and testing on the synthesized actions. Since their model initializes the sequences with ground-truth from the test data, this could positively bias their scores. Therefore, we omit test data from training, in our analysis in Table 1 of the main paper, both for our approach and IMoS. We show however in this supplemental material (Table 7) that our model is still able to outperform IMoS even when all the GRAB data, including test data, is included in training the classifier. This is a setup which is more favorable for IMoS as the first frame of synthesized IMoS motion comes from test data. This re-confirms the ability of our method to encode diverse motion features that represents realistic hand-object interactions.
| Method |
DIV () | MM () | KID () | FID () | |||||
Real Mocap | 1.0 | 1.0 | 1.13580.0164 | 0.31050.0163 | - | - | ||||
| IMoS (Ghosh et al., 2023) | 0.7017 | 0.6754 | 1.10530.0128 | 0.28820.0114 | 0.005811 | 0.6267 | |||||
| DiffHO | 0.8125 | 0.9028 | 1.14400.0115 | 0.31760.0127 | 0.005346 | 0.8342 | |||||
Real Mocap | 1.0 | 1.0 | 1.13240.0119 | 0.23550.0044 | - | - | ||||
| IMoS* (Ghosh et al., 2023) | 0.7161 | 0.8000 | 1.11730.0112 | 0.25130.0104 | 0.006038 | 0.7101 | |||||
| DiffHO | 0.7836 | 0.8125 | 1.12250.0120 | 0.21740.0074 | 0.005494 | 0.8821 |
FID and KID metrics.
It’s been shown by earlier studies (Noguchi and Harada, 2019), (Jayasumana et al., 2023), (Chong and Forsyth, 2020) that Fréchet Inception Distance (FID) metric is biased and not stable for calculating the distances for small datasets. Since the GRAB subject split only contains 144 motions, in addition to FID, we also report the more stable Kernel Inception Distance (KID) metric which provides more reliable scores with fewer samples (Table 8). Table 8 also includes the standard deviations for the multimodality and diversity metrics based upon 20 repetitions of the evaluation.
| Method | DIV () | MM () | KID () | FID () | ||||
Ground-truth | 1.06320.0131 | 0.21550.0091 | - | - | |||
| IMoS (Ghosh et al., 2023) | 1.02370.0158 | 0.25380.0103 | 0.008065 | 0.6267 | ||||
| DiffHO | 1.07570.0129 | 0.30370.0079 | 0.006697 | 0.8342 | ||||
Ground-truth | 1.07570.0129 | 0.20020.0063 | - | - | |||
| IMoS* (Ghosh et al., 2023) | 1.04710.0129 | 0.22020.0093 | 0.008641 | 0.6593 | ||||
| DiffHO | 1.09420.0125 | 0.23070.0060 | 0.007503 | 0.8425 |
Full-sequence training.
Given an initial pose of the grasp moment, IMoS only synthesizes post-grasp sequences. To compare our results to IMoS in Table 1 of the main paper, we evaluated our models only on the post-grasp sequence part of the synthesis. In Table 9, we also provide diversity, multimodality, FID, KID and action recognition scores computed over the full sequence including pre- and post-grasp stages. We compare our results against real motions from ground-truth and demonstrate that we can achieve scores similar to ground-truth, which demonstrates our capability to synthesize realistic hand-object interactions.
| Method |
DIV () | MM () | KID () | FID () | ||||
| Real Mocap (Full-sequence) | 1.0 | 1.0 | 1.09830.0059 | 0.22350.0042 | - | - | ||||
| DiffHO (Full-sequence) | 0.7720 | 0.8031 | 1.09480.0120 | 0.15550.0057 | 0.0067 | 0.7993 |
C.2. Text Input Evaluation Experiments
In Table 10, we provide further details for the text input to the diffusion model. We first compare IMoS and our model trained on the subject split and tested on the object split. Notably, our model outperforms IMOS by 9.2% in action correctness, a significant improvement. When trained and tested on our unseen object splits, we find that our model remains competitive with IMoS. However, we notice that our model struggles with classes like apple and train, where new motions are introduced that differ from those in the training set. Our model must learn to generalize from eating bananas to apples, and while it achieves this, its performance decreases to 81.8%. Similarly, for the train class, our model’s accuracy drops to 50% since the unique motions of playing with a train were not observed during training. Nevertheless, our model demonstrates the ability to generalize to different object sizes, as it generates motions corresponding to recognizable actions for varied sizes of cylinders, pyramids, and torus in the training set.
Finally, we train and test our model with the augmented sentences we proposed for GRAB, leading to improved generalization, especially for the classes “train” and “apple”. These new sentences provide detailed motion descriptions, such as “The person picks up the train using their right hand, then rides it through the air close to the table surface, and finally puts it on the table using their right hand.” The augmented sentences enhance generalization and provide control over specifics like handedness, motion type, and physical location. Testing a model trained solely with baseline sentences on augmented sentences reveals a decrease in generalization, as anticipated. We observed a significant 38.7% performance gap between baseline and augmented sentences, showing the importance of text style. This highlights the limitations of relying solely on simple sentences for training and underscores the value of incorporating diverse and detailed descriptions to enhance model performance and generalization capabilities.
C.3. Detailed Investigation User Study
In our user study, we provide each user with 30 side-by-side comparisons. In our investigation, our model showed low realism but high diversity for the “pass” action, influenced by users perceiving “pass” action label in the GRAB dataset as transferring an object between hands, not to another person. In the “offhand” category, our model excelled in realism, while IMoS exhibited artifacts. This is due to the fact that IMoS determines the hand which is in contact with the object according to ground-truth, which leads to a jump of the object for the “off-hand” category.
C.4. More Qualitative Results and Failure Cases
We provide more qualitative results of our model with simple instructions in Fig. 5(a), for grasp guidance in Fig. 5(b), and with detailed instructions in Fig. 5(c). Furthermore, we present three failure cases: the generated action does not match the text prompt (Fig. 6(a)), the grasp reference is ignored during guidance (Fig. 6(b)), or the hand in the text prompt is ignored (Fig. 6(c)).
| Split | Textual input form |
Action correctness over object classes | ||||||||||||||||||
| Method | Train split | Test split | Train input | Test input | Overall | apple | mug | train |
elephant |
| IMOS* | subject | object | simple | simple | 90.3 | 90.9 | 89.5 | 77.8 | 100 | 80.0 | 100 | 83.3 | 100 | ||||||||
| Ours | subject | object | simple | simple | 99.5 | 100 | 97.4 | 100 | 100 | 100 | 100 | 100 | 100 | ||||||||
| Ours | object | object | simple | simple | 89.7 | 81.8 | 92.1 | 50.0 | 100 | 100 | 100 | 100 | 100 | ||||||||
| Ours | object | object | simple | comprehensive | 54.8 | 40.9 | 55.3 | 16.7 | 52.6 | 80.0 | 75 | 50.0 | 68.8 | ||||||||
| Ours | object | object | comprehensive | comprehensive | 93.5 | 95.5 | 94.7 | 77.8 | 100 | 95.0 | 93.6 | 100 | 93.6 | ||||||||
Appendix D Computational Resources
The training of our model takes approximately 48 GPU hours on a single NVIDIA V100. Our model has a throughput of 32 samples per second. The total inference time for 32 samples is approximately 300 seconds.

### v2

@@
-DiffHO: Diffusion-Based Synthesis of Hand-Object Interactions from Textual Descriptions
+DiffH2O: Diffusion-Based Synthesis of Hand-Object Interactions from Textual Descriptions
 Abstract.
-Generating natural hand-object interactions in 3D is challenging as the resulting hand and object motions are expected to be physically plausible and semantically meaningful.
-Furthermore, generalization to unseen objects is hindered by the limited scale of available hand-object interaction datasets.
-We propose DiffHO, a novel method to synthesize realistic, one or two-handed object interactions from provided text prompts and geometry of the object.
-The method introduces three techniques that enable effective learning from limited data.
-First, we decompose the task into a grasping stage and a text-based interaction stage and use separate diffusion models for each.
-In the grasping stage, the model only generates hand motions, whereas in the interaction phase both hand and object poses are synthesized.
-Second, we propose a compact representation that tightly couples hand and object poses.
-Third, we propose two different guidance schemes to allow more control of the generated motions: grasp guidance and detailed textual guidance.
-Grasp guidance takes a single target grasping pose and guides the diffusion model to reach this grasp at the end of the grasping stage, which provides control over the grasping pose.
-Given a grasping motion from this stage, multiple different actions can be prompted in the interaction phase.
-For textual guidance, we contribute comprehensive text descriptions to the GRAB dataset and show that they enable our method to have more fine-grained control over hand-object interactions.
-Our quantitative and qualitative evaluation demonstrates that the proposed method outperforms baseline methods and leads to natural hand-object motions.
-Moreover, we demonstrate the practicality of our framework by utilizing a hand pose estimate from an off-the-shelf pose estimator for guidance, and then sampling multiple different actions in the interaction stage.
-The project page along with videos can be found at https://diffh2o.github.io/
+We introduce DiffH2O, a new diffusion-based framework for synthesizing realistic, dexterous hand-object interactions from natural language.
+Our model employs a temporal two-stage diffusion process, dividing hand-object motion generation into grasping and interaction stages to enhance generalization to various object shapes and textual prompts.
+To improve generalization to unseen objects and increase output controllability, we propose grasp guidance, which directs the diffusion model towards a target grasp, seamlessly connecting the grasping and interaction stages through a motion imputation mechanism.
+We demonstrate the practical value of grasp guidance using hand poses extracted from images or grasp synthesis methods.
+Additionally, we provide detailed textual descriptions for the GRAB dataset, enabling fine-grained text-based control of the model output.
+Our quantitative and qualitative evaluations show that DiffH2O generates realistic hand-object motions from natural language, generalizes to unseen objects, and significantly outperforms existing methods on a standard benchmark and in perceptual studies.
 1. Introduction
-Modeling and understanding hand-object interactions is an important and crucial task to empower machines to interact with and assist humans.
+Modeling and understanding hand-object interactions (HOI) is an important and crucial task to empower machines to interact with and assist humans.
 Being able to seamlessly generate hand-object motions holds the promise to enable synthetic data generation (Leonardi et al., 2023), assist robots in training human-robot interactions in simulation, or enhance realism in virtual manipulation experiences.
@@
 Generating realistic hand-object interactions in 3D is challenging as the resulting motions are required to be plausible in several different aspects.
-First, the motions must be plausible in terms of geometry, such hand and object intersections are minimized and the grasp appears stable.
+First, the motions must be plausible in terms of geometry, such that hand and object intersections are minimized and the grasp appears stable.
 Second, the motions must be plausible in terms of semantics, such as the hands respect natural object affordances (e.g., a cup is grasped by its handle and not flipped upside down).
 Third, the motions must be plausible in terms of time, such as the hand and object motions are synchronized and the dynamics appear natural.
-Another challenge in generating hand-object interactions comes from the limited scale of existing hand-object datasets, which are around 10x smaller than human motion datasets (Mahmood et al., 2019) and 1000x smaller than image datasets (Deng et al., 2009).
-Several recent works have successfully leveraged diffusion models for full-body human motion generation but either neglect objects (Tevet et al., 2023; Karunratanakul et al., 2023) or focus only on coarse motion of larger objects (e.g., a chair) without finger predictions (Xu et al., 2023).
-Another method, IMoS (Ghosh et al., 2023), can generate hand-object interactions with smaller objects (e.g., a cup or stapler) by first optimizing for upper body motions and then optimizing for object poses in a post-processing step.
-However, IMoS assumes that the object is in hand from the beginning, can handle only objects seen during training, and, as shown in our qualitative evaluation, the optimization often yields artifacts (e.g., hand-object interpenetration or non-plausible contact points).
-In this paper, we aim to learn a model that can generate natural, fine-grained hand-object interactions and generalize to objects unseen during training.
-Towards this goal, we contribute a novel diffusion-based framework, DiffHO, that can generate hand-object interactions based on provided text prompts and geometry of the object.
+Another challenge in generating HOI comes from the limited scale of existing hand-object datasets, which are around 10x smaller than human motion datasets (Mahmood et al., 2019) and 1000x smaller than image datasets (Deng et al., 2009).
+Several recent works have successfully leveraged diffusion models for full-body human motion generation but either neglect objects (Tevet et al., 2023; Karunratanakul et al., 2023) or focus only on coarse motion of larger objects (e.g., a chair) without finger predictions (Xu et al., 2023a).
+Another method, IMoS (Ghosh et al., 2023), generates HOI with smaller objects (e.g., a cup or stapler) by first predicting upper-body motions and then optimizing object poses in post-processing.
+However, IMoS assumes that the object is in-hand from the beginning, handles only objects seen during training, and, as shown in our experiments, often yields artifacts (e.g., hand-object interpenetration or non-plausible contacs).
+In this paper, we aim to learn a model that generates natural, fine-grained hand-object interactions and generalize to objects unseen during training.
+Towards this goal, we contribute a novel diffusion-based framework, DiffH2O, that generates hand-object interactions based on text prompts and geometry of the object.
 We propose several techniques to deal with the challenges of data scarcity and object generalization.
@@
 For example, drinking from a cup can be performed with multiple grasps, and given a grasp, multiple actions are possible.
-We use this simple observation to obtain a grasping and an interaction diffusion model, and introduce an inpainting technique called subsequence guidance to allow continuity between the two outputs.
-Different from previous diffusion frameworks that employ coarse-to-fine (2D trajectory-to-full motion) two-stage generation (Karunratanakul et al., 2023), we split the task temporally.
+We use this simple observation to obtain a grasping and an interaction diffusion model, and introduce an inpainting technique called subsequence imputing to allow continuity between the two outputs.
+In contrast to previous diffusion frameworks that employ coarse-to-fine (trajectory-to-full motion) motion generation (Karunratanakul et al., 2023), we split the task temporally and propose a temporal two-stage diffusion model.
 Pose Representation: We introduce a compact representation that models hands in parametric space.
-To couple the hands to the object, we represent the global hand positions relative to the normalized object pose at the initial frame and further include distance information between the 3D hand joints and the object surface.
-Unlike previous representations utilized in human motion diffusion (Tevet et al., 2023) and hand-object interaction (Christen et al., 2022), we find a non-redundant representation directly in parametric space to perform best.
+To couple the hands to the object, we represent the global hand positions relative to the normalized object position at the initial frame, and further include distance information between the 3D hand joints and the object surface.
+While encoding distances helps reason about local shape, representing coordinates relative to the initial frame allow us to better adjust to unseen object geometry compared to per-frame object-relative pose representation used in existing work (Christen et al., 2022).
 Controllability: We provide two ways to add more control to our diffusion models’ output.
 First, we propose control through fine-grained textual descriptions.
-To this end, we contribute textual descriptions to the GRAB dataset and show that they enable more robustness to unseen test prompts and increase controllability, e.g., by allowing to define the interacting hand and the action to perform through the input prompts.
-Second, in grasp guidance, we leverage a single grasp reference which connects the grasping and interaction phase and use it to guide our two diffusion models at inference time.
-This grasp pose can either be obtained from motion capture or from an image-based pose estimator (Pavlakos et al., 2023) as we show in our experiments Section 5.8.
+To this end, we contribute textual descriptions to the GRAB dataset (Taheri et al., 2020) and show that they enable more robustness to unseen test prompts and increase controllability, e.g., by allowing to prompt the interacting hand and the action to perform.
+Second, when a target grasp reference pose for a given object is available, we propose to leverage it to connect the grasping and interaction phases by guiding our two diffusion models at inference time.
+The grasp reference provides a prior about how and with which hand an object should be grasped.
+It can either be obtained from from an image-based pose estimator (Pavlakos et al., 2024) or grasp synthesis (Zhang et al., 2024a) as we show in our experiments in Section 5.7.
 To demonstrate the effectiveness of our approach, we run perceptual studies and experiments on the GRAB dataset (Taheri et al., 2020).
@@
 Our results indicate that our method generates better results across a variety of physics-based and motion metrics.
-To justify our technical contributions, we compare our final model against human-body diffusion baselines by adapting these methods to hand-object interactions.
-We show that we generate motions of higher quality while enabling more control of the outputs.
+To justify our technical contributions, we compare our final model against human-body diffusion baselines by adapting these methods to HOI.
+We show that we generate motions of higher quality while enabling more control of the outputs reaching feasible poses.
 Furthermore, we demonstrate that detailed textual descriptions increase robustness to unseen texts and offer better control of the diffusion model than training with heuristics-generated text descriptions.
-Decoupling and grasp-guidance brings high practical value to DiffHO.
-We show this with the first application of using an image-based pose/shape estimator (Pavlakos et al., 2023) to guide the diffusion to the grasp, and thereafter sampling multiple different actions through text.
-We prove the utility of our method by largely outperforming the state-of-the-art on a standard benchmark and in perceptual studies.
-Our contributions can be summarized as follows:
-To the best of our knowledge, DiffH2O is the first method that synthesizes hand-object interactions on unseen objects from textual descriptions, unlike previous generative models for human-body motion and seen-object HOI synthesis.
-We propose a two-stage diffusion process that splits generation into grasping and action-based interaction.
-To increase the controllability of outputs and improve generalization to unseen objects, we propose grasp guidance and subsequence guidance that can be applied at inference time to the diffusion process.
-We contribute detailed textual annotations to the GRAB dataset along with experiments showing that these textual descriptions increase robustness and enable fine-grained control of the model output.
-Code and data will be made public upon acceptance.
+Decoupling our method into two stages and grasp guidance brings better generalization and high practical value to DiffH2O through controllability, e.g., for downstream tasks such as animation or generating synthetic data at scale.
+We show this with two applications for grasp guidance: i) using an image-based pose estimator (Pavlakos et al., 2024) and ii) leveraging pre-generated grasp poses (Zhang et al., 2024a) to define target reference grasps, and consequently sampling multiple different actions through text in the interaction stage.
+We contribute the following: (i) To the best of our knowledge, DiffH2O is the first method that synthesizes hand-object interactions on unseen objects from textual descriptions.
+(ii) A temporal two-stage diffusion process that splits generation into grasping and action-based interaction, which improves generalization to different textual prompts.
+(iii) Grasp guidance and subsequence imputing that can be applied at inference to the diffusion process to increase the controllability of outputs and improve generalization to unseen objects.
+(iv) Detailed textual descriptions to the GRAB dataset.
+Code and data will be made public upon acceptance***For code and data, see our project website here: https://diffh2o.github.io.
 2. Related Work
 2.1. Hand-Object Synthesis
-Recent efforts in interaction synthesis have largely been driven by the surge of high-quality human-object interaction datasets (Taheri et al., 2020; Brahmbhatt et al., 2019; Hampali et al., 2020; Brahmbhatt et al., 2020; Chao et al., 2021; Kwon et al., 2021; Fan et al., 2023; Liu et al., 2022).
-Some studies focus on generating coarse full-body object interactions (Zhang et al., 2022a; Wang et al., 2021; Lee and Joo, 2023; Hassan et al., 2023; Luo et al., 2021), such as carrying or moving around boxes.
-FLEX (Tendulkar et al., 2023) trains a hand and body pose prior and later optimizes the priors to achieve diverse, static full-body grasps.
-GOAL (Taheri et al., 2022) and SAGA (Wu et al., 2022) use CVAEs to generate approaching motions for full-body grasps, whereas TOHO (Li et al., 2024) models both approaching and manipulation tasks via neural implicit representations.
+Recent efforts in interaction synthesis have largely been driven by the surge of high-quality human-object interaction datasets (Taheri et al., 2020; Brahmbhatt et al., 2019; Hampali et al., 2020; Brahmbhatt et al., 2020; Chao et al., 2021; Kwon et al., 2021; Fan et al., 2023; Liu et al., 2022; Hasson et al., 2019; Yang et al., 2022; Zhan et al., 2024; Wang et al., 2023).
+Some studies focus on generating coarse full-body object interactions (Zhang et al., 2022a; Wang et al., 2021; Lee and Joo, 2023; Hassan et al., 2023; Luo et al., 2021), such as carrying boxes.
+FLEX (Tendulkar et al., 2023) trains a hand and body pose prior and optimizes the priors to achieve diverse, static full-body grasps.
+GOAL (Taheri et al., 2022) and SAGA (Wu et al., 2022) use CVAEs to generate approaching motions for full-body grasps, whereas TOHO (Li et al., 2024b) models both approaching and manipulation tasks via neural implicit representations.
 Similarly, Braun et al.
-(2024) model the full range of motion, leveraging physics simulation and reinforcement learning.
-In contrast to these works, we focus on fine-grained hand-object interactions and generalization to unseen objects.
-Furthermore, we leverage language prompts as input to generate interaction sequences.
-Among methods that model full-body object interaction, closest to our work is IMoS (Ghosh et al., 2023), a two-stage method to generate hand-object interactions on seen objects based on language commands.
-Starting from a grasping state, they first generate body motions and then optimize for object trajectories using a heuristics-based optimization to model hand-object interactions.
-By contrast, our model directly predicts the hand and object poses, models both approaching and manipulation phases, and generalizes to unseen objects.
+(2024) model the full range of motion, leveraging physics simulation and reinforcement learning (RL).
+In contrast, we focus on fine-grained HOI and generalization to unseen objects.
+Among methods that model full-body object interaction, closest to our work is IMoS (Ghosh et al., 2023), a two-stage method to generate HOI on seen objects based on language prompts.
+Starting from a grasping state, they first generate body motions and then optimize for object trajectories using a heuristics-based optimization to model HOI.
+By contrast, our model directly predicts the hand and object poses, models both approaching and manipulation, and generalizes to unseen objects.
 Another line of research focuses on generation of hand-object interaction sequences in isolation from full-body due to wide applications in VR and the need for a dedicated model for fine-grained details of finger motion.
-One prominent solution is to turn to physical simulation and reinforcement learning (Christen et al., 2022; Mandikal and Grauman, 2021; Garcia-Hernando et al., 2020; Rajeswaran et al., 2018; Qin et al., 2021).
-A number of studies propose to learn dexterous manipulation tasks from full human demonstration data either collected via teleoperation (Rajeswaran et al., 2018) or from videos (Garcia-Hernando et al., 2020; Qin et al., 2021).
+One prominent solution is to turn to physical simulation and RL, e.g., by learning dexterous manipulation tasks in simulation from full human demonstration data either collected via teleoperation (Rajeswaran et al., 2018) or from videos (Garcia-Hernando et al., 2020; Qin et al., 2021).
 Mandikal and Grauman (2021) propose a reward function that incentivizes dexterous robotic hand policies to grasp in the affordance region of objects.
+Li et al.
+(2024a) propose a solution for grasping multiple objects at once.
 She et al.
-(2022) propose a new dynamic state representation to generate dexterous grasps.
-D-Grasp (Christen et al., 2022) leverages reinforcement learning and physics simulation to generate diverse hand-object interactions from sparse reference inputs.
-ArtiGrasp (Zhang et al., 2024) extends this to two-handed grasps and generates articulated object motions.
-To model hand-object interaction sequences, another solution is to rely on purely data-driven frameworks.
-For example, (Zhou et al., 2022; Anonymous, 2023) propose methods that enable denoising hand-poses from noisy sequences of hand-object poses.
+(2022) propose a new dynamic state representation to generate dexterous grasps, whereas other works leverage implicit representation (Karunratanakul et al., 2020) or force-closure (Liu et al., 2021).
+D-Grasp (Christen et al., 2022) and Unidexgrasp (Xu et al., 2023b) leverage RL and physics simulation to generate diverse hand-object interactions from sparse reference inputs.
+ArtiGrasp (Zhang et al., 2024b) extends this to two-handed grasps and generates articulated object motions.
+Unidexgrasp++(Wan et al., 2023) proposes a solution that transfers to vision-based inputs, which may be used for robotic grasping.
+These works focus on achieving stable grasping, whereas we model both the grasping and manipulation of objects.
+To model HOI, another solution is to rely on purely data-driven frameworks.
+For example, (Zhou et al., 2022; Liu and Yi, 2024; Taheri et al., 2024) propose methods that enable denoising hand-poses from noisy sequences of hand-object poses.
 Ye and Liu (2012) predict the local hand pose given full body and object motion.
@@
 2.2. Diffusion for Motion Synthesis
-Diffusion models (Sohl-Dickstein et al., 2015) have gained popularity in many domains, such as image generation (Ho et al., 2020), video generation (Yang et al., 2023), audio synthesis (Kong et al., 2021) or hand reconstruction (Ye et al., 2023b, a).
-More recently, diffusion models have also been adapted to human motion synthesis (Tevet et al., 2023; Zhang et al., 2022b).
-Various improvements have been proposed, such as integrating physics (Yuan et al., 2023), scene-awareness (Huang et al., 2023), increasing efficiency by diffusing in a pre-trained latent space (Chen et al., 2023a), or the composition of multiple actions (Athanasiou et al., 2022).
-To enable more controllability during motion synthesis, a two-stage framework, GMD, that guides the diffusion model towards target objectives at inference time is proposed in Karunratanakul et al.
-(2023).
+Diffusion models (Sohl-Dickstein et al., 2015) have gained popularity in many domains such as image generation (Ho et al., 2020), video generation (Yang et al., 2023), audio synthesis (Kong et al., 2021) or hand reconstruction (Ye et al., 2023b, a).
+They have also been adapted to human motion synthesis (Tevet et al., 2023; Zhang et al., 2022b).
+Various improvements have been proposed by integrating physics (Yuan et al., 2023), scene-awareness (Huang et al., 2023), increasing efficiency by diffusing in a pre-trained latent space (Chen et al., 2023a), or the composition of multiple actions (Athanasiou et al., 2022).
+To enable more controllability during motion synthesis, Karunratanakul et al.
+(2023) propose GMD that guides the diffusion model towards target objectives at inference time.
 However, GMD contains design choices specific to human motion generation, by first generating a 2D root trajectory over the whole sequence, and then generating corresponding full-body poses, which is not directly transferable to HOI.
-While these works focus on human motion in isolation from objects, InterDiff (Xu et al., 2023) generates human-object interactions via diffusion models.
+While these works focus on human motion in isolation from objects, InterDiff (Xu et al., 2023a) generates human-object interactions via diffusion models.
 Their work is improved upon by recent concurrent works (Diller and Dai, 2023; Li et al., 2023; Peng et al., 2023) that leverage contact-based predictions in combination with inference time-guidance to improve the interaction quality.
@@
 Most similar to ours, the concurrent work MACS (Shimada et al., 2023) proposes a diffusion model for hand-object motion synthesis.
-Their focus is on synthesizing interactions for a single object with varying mass, while our focus is and generalization to unseen shapes and interaction synthesis conditioned on text.
-3. Denoising Diffusion Probabilistic Models
-Denoising Diffusion Probabilistic Models (DDPM) (Sohl-Dickstein et al., 2015; Ho et al., 2020) aim to model the probability distribution of a given dataset and can model complex distributions of images (Rombach et al., 2022), videos (Ho et al., 2022) or time-series data such as motion sequences (Tevet et al., 2023).
-The diffusion model involves a forward process, which is a Markov chain and consists of sequentially adding Gaussian noise to the data .
-Each step in the forward noising process results in a distribution, , where is a hyper-parameter and , where is the final number of steps.
-Alternatively, can be expressed in terms of as
-where, .
-The reverse process, also called the denoising process is another Markov chain with learned Gaussian transitions starting at .
-The denoised distribution at step is given by, , where are the model parameters to be learnt, is the mean, and is the variance.
-The process of data generation involves sequentially drawing samples from this distribution.
-The variance is usually set to and which is implemented using an auto-encoder trained to match the mean of forward process posterior which is another Gaussian represented as, , where
-and .
-The auto-encoder could directly estimate the mean or in Eq.
-2, or the noise which is used in the noising process of Eq.
-1.
-When estimating the noise, the denoised mean is obtained as , where is the auto-encoder output.
+Their focus is on synthesizing interactions for a single object with varying mass, while our focus is on generalization to unseen shapes and interaction synthesis conditioned on text.
+3. Preliminary: Diffusion Models
+Denoising Diffusion Probabilistic Models (DDPM) (Sohl-Dickstein et al., 2015) model the probability distribution of a given dataset for images (Rombach et al., 2022), videos (Ho et al., 2022) or motion sequences (Tevet et al., 2023).
+The diffusion model involves a forward process, which is a Markov chain consisting of sequentially adding Gaussian noise to the data, and a reverse process to denoise the data gradually.
+The forward process results in a distribution, , where a clean input is gradually noised to , with as the total number of diffusion steps.
+The reverse process models the probability distribution, , where the noisy input is denoised into , which is modeled via a neural network with parameters .
+The process of data generation involves sequential denoising and noising steps.
+The denoising is modeled in auto-encoder fashion predicting either directly the denoised output, ; the mean, , or the noise, , used in the noising process.
 The auto-encoder is trained to minimize
-where, is either , , or .
-Conditional diffusion models generate output conditioned on various kinds of input such as text (Rombach et al., 2022) and other semantic information (Zhang et al., 2023).
-Such a model is obtained by providing the condition as additional input to the auto-encoder resulting in outputs , or , where c is the conditional input encoding.
-3.1. Guiding the diffusion model
+where is either , , or .
+Depending on the application, a different choice of the model output works better in practice, e.g., directly predicting the output for human motion (Tevet et al., 2023) or noise prediction for images (Rombach et al., 2022) or guidance (Karunratanakul et al., 2023).
+Conditional diffusion models generate output conditioned on various kinds of input such as text (Rombach et al., 2022) or other semantic information (Zhang et al., 2023).
+Such a model is obtained by providing the condition as additional signal, c, to the auto-encoder, resulting in outputs , , or .
 Diffusion models can be guided to generate outputs of the required form.
 Two popular methods to guide the diffusion models are classifier guidance (Dhariwal and Nichol, 2021) and classifier-free guidance (Ho and Salimans, 2022).
-We describe here the classifier-guidance method which is more relevant to our work and refer the reader to (Ho and Salimans, 2022) for details on classifier-free guidance.
-The classifier-guidance method approximates , where c is the guidance, by a Gaussian:
-Intuitively, the scaled gradient of the guidance signal, is used to nudge the denoised mean to generate samples in accordance with the guidance c.
-could represent the class probability which can be implemented using a neural-net (Dhariwal and Nichol, 2021), or could represent a cost term that is optimized (Karunratanakul et al., 2023).
-The classifier guidance method has the advantage that any guidance signal that is differentiable can be used to guide the diffusion and can use pre-trained diffusion models without retraining them.
+We describe the classifier-guidance method here and refer the reader to (Ho and Salimans, 2022) for details on classifier-free guidance.
+The classifier-guidance method approximates , where c is the guidance signal.
+Deriving from this, Dhariwal and Nichol (2021) show that the new sampling process can be expressed as
+where indicates the original mean, is the scaling of the gradients, and is a variance scheduler.
+Here the mean is shifted by the scaled gradient.
+Intuitively, the scaled gradient of the guidance signal, nudges the denoised mean to generate samples in accordance with the guidance c.
+could represent the class probability which can be implemented using a neural-net, or could represent a cost term that is optimized.
+The classifier guidance method has the advantage that it can use pre-trained diffusion models without retraining them and can benefit from any guidance signal that is differentiable to guide the diffusion.
 4. Method
@@
 In other words, we want to model a conditional probability distribution with a motion DDPM.
-The optional term is a differentiable goal function that should be minimized, e.g., the distance to a target grasp.
+The optional term can be any differentiable goal function that should be minimized, e.g., the distance to a target grasp in our case (see Sec.
+4.3).
 See Fig.
@@
 In Sec.
-4.1, we introduce a compact representation that tightly couples the hand and object poses and helps in generating realistic hand-object interactions.
+4.1, we introduce our compact representation that couples the hand and object poses.
 We propose a two-stage generation of hand-object interactions in Sec.
@@
 Lastly, in Sec.
-4.4, we introduce our newly-collected detailed textual annotations for the GRAB dataset that enable more controlled generation of hand-object interactions.
+4.4, we introduce our newly-collected detailed textual annotations for GRAB that enable more controlled generation of HOI.
 4.1. Canonicalized Hand-Object Representation
-Our proposed representation of length N contains information about both hands and an object.
-In particular, the object poses are defined as , where is a sequence of 3D object locations, with , and is a sequence of 3D object orientations, with given in 6D representation (Zhou et al., 2019).
-The hands are represented by: , where denotes the left and denotes the right hand.
+Our proposed canonicalized representation , where is the number of steps in the sequence, contains information about both hands and an object.
+The object poses are defined as , where is a sequence of 3D object locations, with , and is a sequence of 3D object orientations, with in 6D representation (Zhou et al., 2019).
+The hands are represented by: , where denotes the left and the right hand, respectively.
 This representation is based on the parametric MANO hand model (Romero et al., 2017).
 Specifically, the global 3D positions of the hands are given by and the global 3D orientations are represented by .
-To achieve more natural poses, we define the local hand pose in a PCA space of the MANO model with the first 24 components as .
-Lastly, denotes distances between each hand joint (including the fingertips) and its closest point on the object mesh.
-As we show in our ablations (see Tab.
-4) this representation reduces physical artifacts such as interpenetration.
-We first normalize the object positions to be at the origin and then adjust the hands’ positions accordingly.
-Hence, all position components are relative to the object origin at the initial frame.
-Due to the limited data, such a representation performs better than a per-frame object-relative representation (Christen et al., 2022; Zhang et al., 2024) as we show in 5.6.
+To achieve more natural poses, we define the local hand pose in PCA space of MANO with the first 24 components as .
+We further encode signed distances, , between each hand joint and its closest point on the object mesh.
+As we show in ablations (see Tab.
+4), this approximation of a signed-distance field (SDF), coupled with our diffusion model reduces physical artifacts.
+Existing works (Christen et al., 2022) encode hand-object poses with a per-frame object-relative representation, which tightly couples the hands to the training object pose and therefore limits generalization to unseen objects.
+In contrast, we represent hand positions relative to the normalized object position at the frame that marks the boundary between the grasp and interaction stage.
+This representation is less variant to specific training object pose and more correlated with the global poses involved in HOI, which is often shared across different training objects in HOI datasets (Chao et al., 2021; Taheri et al., 2020).
+As a result, it yields better generalization and accuracy for HOI across different (unseen) objects, as we show in ablations in Sec.
+5.6.
 4.2. Two-Stage Hand-Object Generation
-We propose to split the synthesis into grasping and interaction stages to facilitate generalization to unseen objects.
-The grasping stage models the motion from an initial pose to an object grasp.
+We split the synthesis into grasping and interaction stages to facilitate generalization to unseen objects.
+The grasping stage models the motion from an initial pose to object grasp.
 The interaction stage performs an action-based manipulation of the object.
 Our two-phase approach is motivated by the fact that any hand-object interaction involves an approaching phase where one or two hands grasp a static object, followed by an intent-driven interaction with the object.
-By decoupling the action-based manipulation from grasping, we can leverage the entire set of motions in the dataset for training the grasping phase, irrespective of the action that is performed in the interaction phase.
-This insight helps counteract the limited scale of the dataset.
+By decoupling the action-based manipulation from grasping, we can leverage the entire set of motions in the data for training the grasping phase, irrespective of the action that is performed in the interaction phase.
+These insights help counteract the limited scale of the data and improve generalization to different textual inputs.
 As shown in Fig.
-2, our grasping and interaction diffusion models are conditioned on three types of inputs to encode intent, namely, 1) text prompts describing the object and the action to be carried out, 2) object mesh which provides the geometry of the object, and 3) the time step in the denoising process.
-We use CLIP (Radford et al., 2021) embeddings to encode the text prompt denoted by , BPS (Prokudin et al., 2019) to encode the object mesh denoted by and an MLP to encode the time step as in previous works (Karunratanakul et al., 2023; Dhariwal and Nichol, 2021).
+2, our grasping and interaction diffusion models are conditioned on three types of inputs, namely, 1) text prompts describing the object and the action to be carried out, 2) object mesh which provides the geometry of the object, and 3) the time step in the denoising process.
+We use CLIP (Radford et al., 2021) embeddings to encode the text prompt denoted by , BPS (Prokudin et al., 2019) to encode the object mesh denoted by and an MLP to encode the time step as in previous works (Dhariwal and Nichol, 2021).
 We provide more implementation details in supplementary material and describe the grasping and interaction phases below in detail.
 4.2.1. Grasping Stage
-The grasping phase is defined as a sequence containing one or two hands approaching a static object from a rest pose until the object is grasped.
+This stage is defined as a sequence containing one or two hands approaching a static object from a rest pose until the object is grasped.
 As the object is static throughout this phase we do not predict the object motion and only model the hand pose sequence.
-In order to obtain the training data for this phase we use heuristics based on object poses to determine sequence boundaries from the larger action sequences and normalize the sequence lengths by interpolating or downsampling the poses.
+To obtain the training data for this phase we use heuristics to determine sequence boundaries from the larger action sequences.
+In particular, the boundary is defined as the first frame where the object’s lateral and vertical velocity surpass a threshold of 0.01m/s and at least 7 vertices of the hand are in contact with the object .
+We then normalize the sequence lengths by interpolating or downsampling the poses.
 The text description in this phase is a generic sentence (”The person grasps the ¡object¿.”).
-We use the diffusion model that directly predicts the noise (see Section 3) in classifier-free manner (Ho and Salimans, 2022) which is more susceptible to inference-time guidance as shown in GMD (Karunratanakul et al., 2023).
+We use the diffusion model that directly predicts the noise in classifier-free manner (Ho and Salimans, 2022) which is more suitable for inference-time guidance (Karunratanakul et al., 2023), e.g., grasp guidance (Section 4.2.3).
 In other words, our grasping diffusion model training loss is given by
-where indicates the synthesized motion at diffusion step .
+where is the synthesized motion at diffusion step and the final output of the grasping stage after denoising with is .
 4.2.2. Interaction Stage
-This phase comprises the motion of both hands and the object after a grasp has been established.
-In this phase, the hands manipulate and interact with the object according to the action defined through the textual prompt.
-We obtain the training data for this phase in a similar way as the grasping phase based on object pose heuristics.
-We employ a diffusion model in a classifier-free manner (see Section 3), i.e., we directly predict the denoised output.
-This is shown to result in less noisy, higher quality motion than (but less susceptibility to guidance) (Karunratanakul et al., 2023), which is critical during the interaction phase where the hand-object interaction takes place.
+This stage comprises the motion of both hands and the object after a grasp has been established.
+The hands manipulate and interact with the object according to the action defined through the textual prompt.
+We use the same sequence boundary determined in the grasping stage to get the training data for interaction.
+We employ a diffusion model in a classifier-free manner (see Section 3), i.e., we directly predict the denoised output , that result in less noisy, higher quality motion than diffusion model for the interaction phase that does not involve inference-time grasp guidance.
 Our training loss is given by
@@
 Maintaining a seamless transition between the grasping and interaction outputs is critical to generating realistic motion sequences.
-We rely on inference-time guidance (Karunratanakul et al., 2023) and motion imputing on the interaction model to achieve a smooth transition.
-To this end, the interaction diffusion model is trained to generate the complete sequence (including grasping).
-We then impute the entire sequence generated by the (guided) grasping model to the start of the interaction sequence and apply guidance.
-To perform the imputation we define respectively the projections and that resize the grasp reference or sub-sequence poses to that of by filling in zeros.
-Similarly, and represents the imputation regions of the grasp reference and sub-sequence on .
-At each denoising step, the imputated denoised output, , is given by
-where and represents elementwise multiplication. We use the classifier guidance mechanism described in Sec. 3.1 and approximate the gradient signal for sub-sequence guidance as
-where denotes the set of grasping poses and denotes the generated grasping poses from the grasping diffusion model.
-We show in Sec.
-5.6 that this technique leads to a smoother transition between pre- and interaction compared to a single frame transition.
+We utilize motion imputing on the interaction model to achieve a smooth transition.
+In particular, we impute the start of the interaction sequence from the entire sequence obtained by the (guided) grasping model.
+Intuitively, this adjusts the generative process for interaction stage based on the observed grasping motion.
+To perform the imputation, we define the projection that resizes the hand poses from the grasping stage to hand-object poses from the interaction stage by filling in zeros, both temporally and spatially.
+Similarly, represents the imputation regions of the grasp sequence on .
+At each denoising step, the imputed denoised output is given by
+where is elementwise multiplication. Our proposed technique, dubbed subsequence imputing, leads to smoother transition between grasp interaction compared to single frame transition (see Sec. 5.6).
 4.3. Grasp Guidance
-To add controllability to the generation and improve generalization to unseen objects, we introduce grasp guidance for the grasping phase, which can be optionally be added at inference time.
-We leverage a single grasp frame, defined as the last frame of the grasping phase, that provides a prior about how and with which hand an object should roughly be grasped.
-Such a grasp can be obtained from motion capture or from an off-the-shelf hand-object pose estimator (Pavlakos et al., 2023) as demonstrated in Sec.
-5.8.
-We provide inference-time guidance to our trained diffusion model to grasp the object in a pre-defined manner.
-More specifically, we approximate the gradient of the guidance signal in Eq.
+To add controllability to the generation and improve generalization to unseen objects, we introduce grasp guidance for the grasping stage, which can be optionally added at inference time via the goal function (see Sec.
+3 and Eq.
+2).
+We leverage a single grasp frame as target grasp (hand pose) for the last frame of the grasping stage.
+It provides a prior about how and with which hand an object should be grasped.
+Such a grasp can be obtained from grasp generation (Zhang et al., 2024a) or from an off-the-shelf hand-object pose estimator (Pavlakos et al., 2024) as demonstrated in Sec.
+5.7.
+We aim to minimize the goal function and approximate the gradient of the guidance signal in Eq.
 2 as
-where is the hand(s) pose in the last frame of the grasping sequence, and is the pre-defined grasp pose.
-| Method | Backbone | SD [] () | OD []() | IV [] () | ID [] () | CR | MM () | DIV () |
-Real Mocap (GRAB) | - | - | 0.183 | 5.65 | 5.89 | 0.08 | 0.22 | 1.06 | 0.708 | 0.729 | ||||
-| IMoS (Ghosh et al., 2023) | CVAE | 0.002 | 0.149 | 7.14 | 11.47 | 0.05 | 0.25 | 1.02 | 0.579 | 0.588 | |||||
-| DiffHO | Transformer | 0.088 | 0.185 | 6.65 | 8.39 | 0.067 | 0.33 | 1.09 | 0.760 | 0.810 | |||||
-| DiffHO | UNet | 0.109 | 0.188 | 6.02 | 7.92 | 0.064 | 0.30 | 1.08 | 0.833 | 0.875 | |||||
-Real Mocap (GRAB) | - | - | 0.183 | 3.40 | 5.63 | 0.071 | 0.20 | 1.08 | 0.683 | 0.784 | ||||
-| IMoS* (Ghosh et al., 2023) | CVAE | 0.002 | 0.132 | 10.38 | 12.45 | 0.048 | 0.22 | 1.05 | 0.561 | 0.581 | |||||
-| DiffHO | Transformer | 0.133 | 0.185 | 7.99 | 10.87 | 0.073 | 0.24 | 1.09 | 0.750 | 0.803 | |||||
-| DiffHO | UNet | 0.134 | 0.179 | 9.03 | 11.39 | 0.086 | 0.23 | 1.09 | 0.755 | 0.837 |
-4.4. Textual Augmentations
+where is the hand(s) pose in the last frame of the grasping sequence after denoising and the cost term minimizes the distance of the hand(s) pose to the target grasp.
+Depending on the availability, the guidance signal may include one or two hands.
+Since the diffusion process is a denoising model of a motion sequence, it relates individual frames to other frames in the sequence.
+Hence, by backpropagating through the diffusion model (via autodiff), we can compute a dense guidance signal for all frames, even if only sparse guidance signals (one in our case) are provided.
+4.4. Detailed Text Descriptions
 Our model generates hand-object interactions based on textual inputs.
-To the best of our knowledge, there is currently no hand-object interaction dataset that contains detailed textual descriptions.
+To the best of our knowledge, there is currently no HOI dataset that contains detailed textual descriptions.
 For instance, the GRAB dataset (Taheri et al., 2020) only provides categorical action labels.
-To address this limitation, we auto-generated sentences using the template “the person ¡verb¿ + ¡object¿” (Ghosh et al., 2023) (referred to as “simple” text).
+To address this, we auto-generated sentences using the template “the person ¡verb¿ + ¡object¿” (Ghosh et al., 2023) (referred to as “simple” text).
 However, these sentences lack detailed information, which limits the controllability of the model via textual inputs.
-Therefore, we contribute carefully annotated textual descriptions of the GRAB dataset (referred to as “detailed” sentences).
+Therefore, we contribute carefully annotated textual descriptions of GRAB (referred to as “detailed” sentences).
 We instructed our annotators to carefully watch each ground truth video and describe the actions in three distinct stages of pre-action, action, and post-action.
 These descriptions are further augmented with hand and positional information.
-An example from the dataset is as follows: ”The person picks up the apple with the right hand, passes it to their left hand, and then places it back with their right hand”.
-These detailed descriptions enhance the dataset, enabling more accurate evaluation and precise control of hand-object interactions as we show in Sec.
-5.7.
+An example from the dataset is as follows: ”The person picks up the apple with the right hand, passes it to their left hand, and places it back with their right hand”.
+These detailed descriptions enhance the dataset, enabling more accurate generation and precise control of HOI as we show in Sec.
+5.8.
+| Method | Backbone | SD [] () | OD []() | IV [] () | ID [] () | CR |
+| Unseen subject split | IMoS (Ghosh et al., 2023) | CVAE | 0.002 | 0.149 | 7.14 | 11.47 | 0.05 | 0.588 | |
+| DiffH2O | Transformer | 0.088 | 0.185 | 6.65 | 8.39 | 0.067 | 0.810 | ||
+| DiffH2O | UNet | 0.109 | 0.188 | 6.02 | 7.92 | 0.064 | 0.875 | ||
+| Unseen object split | IMoS* (Ghosh et al., 2023) | CVAE | 0.002 | 0.132 | 10.38 | 12.45 | 0.048 | 0.581 | |
+| DiffH2O | Transformer | 0.133 | 0.185 | 7.99 | 10.87 | 0.073 | 0.803 | ||
+| DiffH2O | UNet | 0.134 | 0.179 | 9.03 | 11.39 | 0.086 | 0.837 |
 5. Experiments and Results
@@
 Our method focuses on synthesizing detailed hand-object motion based on text input.
-To the best of our knowledge, DiffHO is the first method to tackle this problem.
+To the best of our knowledge, DiffH2O is the first method to tackle this problem.
 While there is no direct baseline for hand-object motion synthesis based on language, the closest to our approach is IMoS (Ghosh et al., 2023), which focuses on text-based whole-body human-object interaction synthesis.
@@
 Since our paper focuses on hand articulations and object motions, we omit the full-body movements from our evaluations.
-To compare our method against human-body diffusion models, we adopted these methods to the HOI setting.
-All variants use our proposed representation.
-MDM (Tevet et al., 2023) does not use guidance, and for GMD (Karunratanakul et al., 2023) we use gradients to guide the model towards the grasp frame, but omit the 2D-trajectory generation stage because it is specific to human motions.
+To further compare DiffH2O against human-body diffusion models, such as (Tevet et al., 2023) and (Karunratanakul et al., 2023), we adopted them to the HOI setting.
+All variants use our proposed pose representation.
+MDM (Tevet et al., 2023) does not use guidance, and for GMD (Karunratanakul et al., 2023), we use gradients to guide the model towards the grasp frame, but omit the 2D-trajectory generation stage because it is specific to human motions.
 5.2. Data
-GRAB: We utilize the subject-based split of the GRAB dataset (Taheri et al., 2020) proposed in IMoS to run a direct comparison.
-However, as this split does not contain unseen objects, we also create a new split featuring unseen objects based on the criteria of object similarity and semantics (see supplemental material for details).
-HO-3D: To further evaluate guidance and object generalization, we estimate hand-object poses on the HO-3D dataset (Hampali et al., 2020) and guide our model to an estimated target pose.
+5.2.1. GRAB
+We utilize the subject-based split of the GRAB dataset (Taheri et al., 2020), which contains 1335 sequences, proposed in IMoS to run a direct comparison.
+However, as this split does not contain unseen objects, we also create a new split featuring unseen objects based on object similarity and semantics (see supplemental material).
+To assess our model’s ability to generalize to unseen classes, we introduce a new split comprising unseen objects that were not present in training.
+Our train and test split contains 1125 and 210 sequences, respectively, which we mirror to a total of 2250 and 420 sequences.
+To split the GRAB sequences into the grasping and interaction stages for training, we apply the heuristics described in Section 4.2.1 Our test set is composed of the following objects: “apple, mug, train, elephant, alarm clock, small pyramid, medium cylinder and large torus”.
+Note that the training set includes pyramids, cylinders, and torus in other sizes.
+We exclude small pyramids, medium cylinders, and large torus from the training set to test our models’ ability to generalize to different sizes of the same objects To evaluate grasp guidance in a controlled manner, we utilize the frame of the unseen test split that indicates the transition between the grasp and interaction stage.
+Note that the grasp reference contains both hands in this case, even if one hand is not grasping the object.
+5.2.2. HO-3D
+To further evaluate guidance and object generalization, we estimate poses on the HO-3D dataset (Hampali et al., 2020) and guide our model to an estimated target hand pose.
 5.3. Evaluation Metrics
-Following previous works on motion synthesis (Ghosh et al., 2023; Tendulkar et al., 2023; Taheri et al., 2022; Braun et al., 2024), we report several physics-based metrics and evaluate our model’s accuracy through metrics that assess motion diversity and action features.
-Physics Based Metrics: We report the interpenetration volume (IV) as the amount of MANO vertices that penetrate the object mesh and the maximum interpenetration depth (ID).
-Furthermore, we compute the contact ratio (CR) as the average ratio of hand vertices that are in contact, which is defined as points within a 5mm threshold of the object mesh.
-Motion Diversity Metrics: We report the sample diversity (SD) between wrist trajectories in motion space.
-We sample the same input conditions five times and compute the pairwise Euclidean distance between the samples.
-We then report the mean sample diversity over all test prompts.
+We report several physics-based metrics and motion diversity metrics following previous works (Ghosh et al., 2023; Tendulkar et al., 2023; Taheri et al., 2022; Braun et al., 2024).
+Please see supplementary material for additional metrics based on action features.
+5.3.1. Physics Based Metrics
+We report the interpenetration volume (IV) as the number of MANO vertices penetrating the object mesh and the maximum interpenetration depth (ID).
+Additionally, we compute the contact ratio (CR) as the average ratio of hand vertices within 5mm of the object mesh.
+5.3.2. Motion Metrics
+We report the sample diversity (SD) between wrist trajectories in motion space.
+We sample the same input conditions 5 times, compute the pairwise Euclidean distance between the samples and report the mean SD over all test prompts.
 We also provide the overall diversity (OD), which measures the pairwise Euclidean distance between all test samples.
-Since the wrist positions are normalized with respect to the initial object pose, no further normalization is applied.
-Guidance Metrics: To assess how well the models adhere to the guidance signal, we report the mean error between the predictions and the reference grasp (GE), the ratio at which the diffusion model maintains the correct handedness according to the reference (HA).
+Furthermore, we report the action recognition accuracy (AR) as in (Ghosh et al., 2023).
+5.3.3. Guidance Metrics
+To assess how well the models adhere to the guidance signal, we report the mean error between the predictions and the reference grasp (GE), and the ratio at which the diffusion model maintains the correct handedness according to the reference (HA).
 Furthermore, we evaluate the wrist velocities at the transition between grasping and interaction phase ().
-Action Feature Based Metrics: To assess the naturalness and variety of the generated motions, we use the action recognition accuracy, diversity, and multimodality metrics, following (Ghosh et al., 2023).
-We provide additional metrics (e.g.
-FID, KID), standard deviations and details about how these metrics are computed in our supplemental material.
-5.4. Comparison to the State-of-the-Art
-We conduct a set of experiments to compare to relevant baselines quantitatively.
-To demonstrate the generality of our method, we present qualitative examples of our method in Fig.
-3 and supplementary materials (along with failure cases).
-Setting 1 - Interaction Only: We compare our approach against the state-of-the-art human-object motion synthesis approach of IMoS in their postgrasp-only setting and report results in Tab.
-1.
-To ensure a fair comparison in this experiment, no grasping and neither grasp guidance nor our detailed text annotations are used.
-We test our hand-object representation with two backbones, MDM’s transformer (Tevet et al., 2023) and GMD’s UNet (Karunratanakul et al., 2023).
-Our model significantly outperforms IMoS across all physics and motion diversity metrics as well as on the action recognition accuracies.
-In particular, our method yields motions that are significantly more diverse (higher SD and OD), exhibit less interpenetration (lower IV and ID), and encode motion features that align with action type (higher action recognition accuracies).
-Note that IMoS assumes the availability of a ground-truth test motion to initialize the first few frames with it, while ours predicts the entire sequence.
-Yet, our approach achieves considerable performance improvements.
-While IMoS provides compelling realistic full body synthesis results, fine-grained finger motions as well as post-optimized object motions lag behind the realism and diversity demonstrated by our method, as shown in Fig.
-3 and in our supplementary video.
-The higher scores of DiffHO compared to real mocap from the GRAB dataset in Tab.
-1 can occur because the action classifier is trained on GRAB’s training-split, which is the distribution our model covers,whereas some of the test sequences in GRAB (real mocap) have quite distinct motions that the classifier may misclassify.
-However, the high scores demonstrate the representative power of DiffHO for generating realistic hand-object sequences that capture inherent motion and action semantics.
-Our method with a UNet backbone outperforms a transformer backbone in most metrics, which is in line with GMD’s findings that UNet works better in low-data regimes.
-To assess the action recognition accuracy, we train two models, one that uses only hand 3D joints as input and another incorporating object 3D position as additional input.
-We demonstrate that DiffHO’s synthesized object motions contribute to larger performance improvements compared to the post-optimized motions computed by IMoS ( improvement for ours vs for IMoS) as shown in Tab.
-1.
-This confirms the benefits of jointly modeling hand-object motion instead of decoupling it into separate components.
 | Method | IV [] () | GE | HA | |
 | MDM | 8.98 | 0.38 | 0.45 | 0.14 |
-| GMD | 9.03 | 0.49 | 0.46 | 12.68 |
+| GMD | 7.47 | 0.30 | 0.58 | 0.20 |
 | Ours | 7.40 | 0.12 | 0.87 | 0.23 |
-Setting 2 - Grasping + Interaction: In this experiment, we switch to the setting of generating the whole sequence of grasping and interaction, and are interested how well the models can be guided while generating high quality motions (without physical artifacts).
-We report results in Tab.
-2 and find that our method outperforms baselines from human motion diffusion.
-In particular, our method better adheres to the grasp frame in terms of grasp error (GE) and handedness accuracy (HA), and achieves more natural grasps (lower IV).
-This justifies our contributions of two-stage separation and subsequence guidance over previous diffusion models.
-5.5. Perceptual Study
+5.4. Comparison to the State-of-the-Art
+We conduct a set of experiments to compare to relevant baselines quantitatively. We present qualitative examples of our method in Fig. 4 and failure cases in Fig. 5.
+5.4.1. Setting 1 - Interaction Only
+We compare our approach against the closest existing work, IMoS, in their interaction-only setting and report results in Tab.
+1.
+To ensure a fair comparison with IMoS, we evaluate our model without the grasp guidance and without our detailed text annotations.
+We test DiffH2O with two backbones, transformer as in (Tevet et al., 2023) and UNet as in (Karunratanakul et al., 2023).
+Our model significantly outperforms IMoS across all physics and motion diversity metrics.
+In particular, our method yields motions that are significantly more diverse (higher SD and OD), exhibit less interpenetration (lower IV and ID), and align better with the action type (higher AR).
+Note that IMoS assumes the availability of a ground-truth test motion to initialize the first few frames, while ours predicts the entire sequence.
+Yet, our approach achieves considerable performance improvements.
+While IMoS provides compelling realistic full body synthesis results, fine-grained finger motions and object motions lag behind the realism and diversity demonstrated by our method, as shown in Fig.
+3 and in our supplementary video.
+In Section 5.5, we further provide a perceptual study to compare our method against the baseline.
+5.4.2. Setting 2 - Grasping + Interaction
+In this experiment, we focus on generating entire sequences of grasping and interaction, evaluating how well the models can produce high-quality motions without physical artifacts and adhere to guidance signals.
+Results in Tab.
+2 show our method outperforms human motion diffusion baselines (MDM/GMD) with better grasp accuracy (GE), handedness (HA), and naturalness (IV).
+This validates our temporal two-stage design with grasp guidance and subsequence imputing over previous diffusion models for the setting of hand-object interactions.
+5.5. Perceptual User Study
 To evaluate the visual quality of our motions, we conducted a perceptual user study on a set of 21 participants, who were randomly assigned one of four sets, each consisting of 30 randomly selected sequences.
@@
 We asked the following questions to the participants: “Which sequence is more realistic?” and “Which sequence has more more variety in motion?”.
-We defined a sequence to be more realistic if the overall motion looks human-like, e.g., if the objects are grasped realistically and there are less artifacts such as floating objects or interpenetration.
+We defined a sequence to be more realistic if the overall motion looks human-like, e.g., if the object is grasped realistically and there are less artifacts such as floating objects or interpenetration.
 Moreover, a sequence has more variety if the object is manipulated multiple times and in distinct ways.
@@
 We obtained statistically significant Fleiss’ kappa scores (Fleiss, 1971) of 0.34 and 0.43 for Realism and Diversity, respectively, at p¡0.05.
-Please see supplemental material for more details.
+In our investigation, our model showed low realism but high diversity for the “pass” action, influenced by users perceiving “pass” action label in the GRAB dataset as transferring an object between hands, not to another person.
+In the “offhand” action category of GRAB, where an object is passed from one hand to another (see Figure 3 and video 02:51), our model excelled in realism, while IMoS exhibited artifacts.
+This is due to the fact that IMoS determines the hand which is in contact with the object according to ground-truth, which leads to a jump of the object for the “off-hand” category.
 5.6. Ablation Studies
@@
 To evaluate the controllability gained through guidance, we compare our final method with several variants in Tab.
-3.
+3 for the full setting that generates the whole sequence with grasping and interaction (Setting 2).
 We use a base model which indicates a single stage model without guidance.
-We then gradually add our contributions, first the grasping and interaction task separation (2-ST), the grasp guidance (GG) and the sub-sequence guidance in the interaction phase (SS) as described in Sec.
+We then gradually add our model components: (i) 2-stage design that decouples grasping and interaction (2-ST), (ii) grasp guidance (GG), and, (iii) subsequence imputing in the interaction phase (SI) as described in Sec.
 4.2.
@@
 Adding guidance allows for better performance in grasp error and hand accuracy, however, a high indicates a sudden jump of the hands between grasping and interaction.
-To mitigate this, we introduce the smooth transition technique that leads to the best overall score at the cost of a slight increase in grasp error and decrease in hand correctness.
-The interpenetration volume is lowest with our final model, indicating the best generalization to unseen objects.
-To validate our canonicalized representation, we compare against a variant inspired by D-Grasp (Christen et al., 2022), our representation without SDF, using object-relative positions at every step, and finally our representation of object-relative with regards to the initial frame.
-Our final representation achieves less interpenetration and better contact modeling.
+Our subsequence imputing mitigates this and leads to the best overall score at the cost of a slight increase in grasp error and decrease in hand correctness.
+The IV is lowest with our final model, indicating the best generalization to unseen objects.
+To validate our representation, we further compare our approach, which normalizes hands with respect to the object position at first frame, against a number of baselines: (i) a variant inspired by D-Grasp (Christen et al., 2022) that involves a redundant representation involving joint angles in Euler space and angular velocities, (ii) our representation without SDF, and, (iii) our representation with per-frame object relative poses.
+As shown in Tab.
+4, our final representation achieves the best overall scores and trade-off with low interpenetration and high contact ratios.
+As discussed in Sec.
+4.1, our representation relative to initial frame is less variant with specific object shapes and more correlated with motion involved in hand-object interaction compared to per-frame object-relative representation.
+This effectively improves generalization for HOI across different objects.
+In addition, SDF feature further improves accuracy by reasoning about local shape.
+The D-Grasp representation has the lowest interpenetration scores, but as indicated by the contact ratio and our qualitative results in the video, the hands are often disconnected from the objects, potentially due its challenge in optimizing in high-dimensional, redundant feature space.
+Additionally, our PCA space helps reduce unnatural hand poses that occur when predicting joints in angle space directly (see supplemental video).
+While the PCA space is beneficial in data-sparse settings, it may hinder more fine-grained finger manipulations compared to joint angle predictions.
 | Component | Metrics | |||||
-| 2-ST | KF | SS | IV [] () | KE | HA | |
+| 2-ST | GG | SI | IV [] () | GE | HA | |
 | 7.48 | 0.31 | 0.52 | 0.21 | |||
@@
 | ✓ | ✓ | ✓ | 7.40 | 0.12 | 0.87 | 0.23 |
-| Pose Representation | Metrics | ||
-| IV [] () | ID | CR | |
-| D-Grasp (Christen et al., 2022) | 5.56 | 8.60 | 0.055 |
-| Ours w/o SDF | 10.79 | 11.42 | 0.073 |
-| Ours w/ object-relative pose | 9.38 | 11.41 | 0.077 |
-| Ours w/ first frame relative pose (Ours) | 9.03 | 11.39 | 0.086 |
+5.7. Generation from Estimated and Pre-Synthesized Pose
+We show two qualitative applications of DiffH2O in video.
+First, we run an image-based pose estimator (Pavlakos et al., 2024) on single images from HO3D (Hampali et al., 2020) to obtain object-relative hand poses.
+Second, we generate grasping poses using GraspXL (Zhang et al., 2024a).
+We then pass these hand poses through our method as reference grasps in grasp guidance (see Sec.
+4.3), together with multiple different textual descriptions.
+This highlights the practicality of our framework given a single grasp reference or grasp motion, to generate multiple diverse sequences.
+See supplementary material for details about how we obtain grasp references and further discussions on their benefits and limitations.
 5.8. Controllability via Text Prompts
-In this experiment, we train DiffHO once with our detailed text labels and once with simple text prompts generated from the GRAB objects and action labels.
-As shown in Tab.
-5, our textual descriptions offer significant advantage by enabling control over handedness.
-To gauge controllability, we measure whether the active hand in the generated motion corresponds to prompted hand.
-Our model, trained with detailed text annotations achieves a correct overall handedness accuracy of 86.5% against 59.3% accuracy when trained with simple text prompts.
-We see a performance drop when training on detailed texts and testing on simple texts (0.731) and vice-versa, which was similarly observed in (Shvetsova et al., 2023) due to a style gap.
-Similarly, we measure action correctness by evaluating the match ratio between the prompted action and the action demonstrated in the generated motion.
-Using detailed textual descriptions enables our model to generate motions that align more accurately with the text input, demonstrated by a 93.5% accuracy compared to the 54.8% accuracy achieved by our model tested with simple text descriptions.
+In this experiment, we train DiffH2O once with our detailed text descriptions and once with simple text prompts generated from the GRAB objects and action labels.
+We measure whether the active hand in the generated motion corresponds to the prompted hand to gauge controllability.
+Results in Tab.
+5 show that detailed text descriptions significantly improve control over handedness, achieving an accuracy of vs.
+with simple prompts.
+Performance drop when training on simple texts and testing on detailed texts (0.702) and vice-versa, which was similarly observed in (Shvetsova et al., 2023) due to a style gap.
+We also measure action correctness by evaluating action recognition accuracy (AR, see Sec.
+5.3).
+Using detailed descriptions enables our model to generate motions that align more accurately with the text input, demonstrated by a 0.887 action accuracy compared to 0.516 with our model tested with simple prompts.
 Since the test inputs are much more diverse with the detailed textual descriptions, this furthermore demonstrates the robustness to unseen sentences.
 We also report the cosine similarity between the different sets of texts as a reference.
-| Train | Test | Action | Hand correctness | Cosine | |||||
-| Input | Input | Correctness | Right | Left | Both | Total | Similarity | ||
+| Pose Representation (unseen objects) | Metrics | ||
+| CR | IV [] | ID | |
+| D-Grasp (Christen et al., 2022) | 0.055 | 5.56 | 8.60 |
+| Ours w/o SDF | 0.073 | 10.79 | 11.42 |
+| Ours w/ object-relative pose | 0.077 | 9.38 | 11.41 |
+| Ours w/ first frame relative pos. (Ours) | 0.086 | 9.03 | 11.39 |
+| Train | Test | AR | Hand correctness | Cosine | |||||
+| Input | Input | Right | Left | Both | Total | Similarity | |||
 0.837 | n/a | n/a | n/a | n/a | 0.72 | ||
@@
 0.887 | 0.869 | 0.862 | 0.75 | 0.865 | 0.66 | ||
-5.8. Motion Generation from Estimated Pose
-We run an image-based pose estimator (Pavlakos et al., 2023) on single images from HO3D (Hampali et al., 2020) to obtain object-relative hand poses.
-We then use these hand poses as reference grasps and pass them through our diffusion framework together with multiple different textual descriptions.
-This highlights the practicality of our framework-given a single grasp reference, we can generate multiple diverse sequences.
 6. Discussion and Conclusion
-In this work, we have introduced a framework to generate plausible hand-object interactions from textual descriptions.
-Specifically, we have proposed a two-stage diffusion method that separates sequences into a grasping and interaction stage and uses single keyframe grasps as guidance to the diffusion model to improve generalization across unseen objects.
-Through our compact representation, hands are coupled to object motions resulting in realistic interactions.
-While we provide a first step towards object generalization, there are limitations of our framework.
+We have introduced a framework to generate plausible hand-object interactions from textual descriptions.
+Specifically, we have proposed a two-stage diffusion framework that decouples sequences into grasping and interaction, and can use single grasps as guidance to the diffusion model to improve generalization to unseen objects.
+While we provide a first step towards HOI synthesis generalization, there are limitations of our framework.
 To remove physical artifacts, physics could be explicitly integrated into the diffusion process (Yuan et al., 2023).
-The inference time is still relatively slow in diffusion models.
+The inference time is still relatively slow.
 To increase efficiency, operating in latent space is an interesting direction to explore (Chen et al., 2023a).
-Lastly, grasp references for guidance which are far away from the training distribution tend to be ignored by our model.
-Therefore, exploring new ways to guide the diffusion model towards out-of-distribution samples is a crucial future research direction.
-Appendix - Supplementary Material
 Appendix A Implementation Details
-DiffHO architecture is based on UNet with Adaptive Group Normalization (AdaGN) which was originally proposed in (Dhariwal and Nichol, 2021) and also adapted in (Karunratanakul et al., 2023) for sequence prediction tasks.
-We adapt this network architecture for our two-stage design, that involves pre-grasp and post-grasp diffusion models.
+While we experiment with a transformer backbone as proposed in MDM (Tevet et al., 2023), DiffH2O’s main architecture is based on UNet with Adaptive Group Normalization (AdaGN) which was originally proposed in (Dhariwal and Nichol, 2021) and also adapted in (Karunratanakul et al., 2023) for sequence prediction tasks.
+We adapt this network architecture for our two-stage design, that involves grasping and interaction diffusion models.
 We provide the hyperparameter settings used in our architecture in Table 6.
 Appendix B Experimental Details
-B.1. Training the action classifier.
-We train a standard RNN action recognition classifier on the GRAB dataset and use the final layer of the classifier as the motion feature extractor for calculating action recognition accuracies as well as diversity (DIV) and multimodality (MM) scores, following (Ghosh et al., 2023) and (Guo et al., 2020).
-We employ the online available code from IMoS and use the same network architecture and parameters to train our model.
-B.2. Unseen object split
-To assess our model’s ability to generalize to unseen classes, we also introduce a new split comprising unseen objects that were not present during training.
-Our test set is composed of the following objects: “apple, mug, train, elephant, alarm clock, small pyramid, medium cylinder and large torus”.
-Note that the training set includes pyramids, cylinders, and torus in other sizes.
-We exclude small pyramids, medium cylinders, and large torus from the training set to test our models’ ability to generalize to different sizes of the same objects
-B.3. Details of the Action Recognition Evaluation Metrics.
-Recognition Accuracy.
-Given a pretrained classifier, we input motions generated by DiffHO as well as our baselines and report top-1 percentage accuracy.
+B.1. Obtaining Grasp References
+As mentioned in the main part of the paper, we show three different ways grasp references can be obtained, namely via existing datasets, pose estimation, or grasp synthesis.
+Note that these procedures for acquiring hand pose references have already been established in related works (Christen et al., 2022; Xu et al., 2023b), albeit for the sole purpose of grasping and not interaction.
+Hand-Object Datasets
+The most straightforward way to attain grasp references is to make use of existing hand-object datasets (see Sec.
+2).
+There are some single frame datasets like ObMan (Hasson et al., 2019) or AffordPose (Ye et al., 2023b) that could be directly used as grasp references.
+However, most existing datasets contain sequential data, hence, we need to identify which frames are potential candidates for hand-object grasps.
+In our paper, we run guidance experiments on our test split of the GRAB dataset (Taheri et al., 2020).
+To acquire suitable frames for guidance, we apply the heuristics described in Sec.
+4.2.1.
+In this case, we can directly use the preprocessed data from GRAB as single keyframes for guidance.
+Image-Based Pose Estimation
+For this experiment, we need to extract the grasp references from images.
+We run an image-based pose estimator (Pavlakos et al., 2024) on single images from HO3D (Hampali et al., 2020) to obtain object-relative hand poses.
+In this experiment, we focus on extracting a single hand pose from the image.
+Please note that we assume the object mesh to be known and at a normalized position to be suitable for our experiments.
+Given the object-relative hand pose, we need to make three modifications to convert it to the same input space our model has.
+First, we have to compute the features for our approximated signed-distance space by finding the closest vertex on the object mesh for each joint position.
+Second, we convert the hand pose from joint angle space to MANO’s PCA space.
+Lastly, because the GRAB dataset is normalized to approach objects from one direction, we have to adjust the orientation of the hand and object to roughly be in line with this direction.
+This could be alleviated by adding data augmentation to the training data in the future.
+We also leave exploring object mesh reconstruction for future work, for example, by utilizing methods that reconstruct the object mesh along with the hand pose estimation (Fan et al., 2024).
+Grasp Synthesis
+In this experiment, we use a grasp synthesis method, GraspXL (Zhang et al., 2024a), to obtain grasp references for guidance.
+To this end, we run GraspXL on several objects to obtain grasping motions.
+As we only require one frame for the guidance, we take the last frame of the grasping motion as the keyframe.
+Since GraspXL allows conditioning on approaching direction, we select grasps whose approaching direction aligns with the GRAB dataset.
+Again, we compute the approximated signed distance by finding the closest vertex on the object mesh for each joint position and convert the hand pose from joint angle to MANO PCA space.
+There is no further conversion needed, as the object poses are already normalized.
+Similar to the experiment with hand poses from images, we only acquire an object-relative grasp and its respective features for one hand.
+B.2. Additional Motion Feature Evaluation Metrics.
+For completion, we also provide metrics based on an action recognition classifier following previous motion synthesis works (Ghosh et al., 2023; Tevet et al., 2023).
+Action Recognition Accuracy (AR).
+Given a pretrained classifier, we input motions generated by DiffH2O as well as our baselines and report top-1 percentage accuracy.
 The recognition accuracy indicates the correlation of the action type and the motion.
-Diversity.
+Diversity (DIV).
 We compute the diversity score by computing the variance of the features extracted from the action classifier across all action categories.
@@
 With , the diversity is defined as
-Multimodality.
-Different from the diversity metric, multimodality computes variance only within a specific action. An overall score is attained based on averaging the variances across all action types.
+Multimodality (MM).
+Different from diversity, multimodality computes variance only within a specific action. An overall score is attained based on averaging the variances across all action types.
 Fréchet Inception Distance (FID).
@@
 It has the assumption that the features computed from the real data and synthesized data have a Gaussian distribution.
-Earlier work demonstrated that for small sample sizes, FID metric is biased and not stable.
+Earlier work demonstrated that for small sample sizes, FID metric is biased and not stable (Noguchi and Harada, 2019).
 Kernel Inception Distance (KID).
@@
 This metric has been found to be more robust when the sample size is small (Noguchi and Harada, 2019).
-| Pre-Grasp | Post-grasp | |
+B.3. Metrics - Training the Action Classifier.
+To provide motion feature metrics (see Sec.
+B.2) based on an action classifier, we train a standard RNN action recognition classifier on the GRAB dataset.
+We use the final layer of the classifier as the motion feature extractor for calculating action recognition accuracies as well as diversity (DIV) and multimodality (MM) scores, following (Ghosh et al., 2023) and (Guo et al., 2020).
+We employ the online available code from IMoS and use the same network architecture and parameters to train our model.
+| Grasping | Interaction | |
 | Parameter | Model | Model |
@@
 | Model avg. beta | 0.9999 | |
-B.4. Visualization Pipeline.
-We use AITViewer (Kaufmann et al., 2022) to visualize motions synthesized by DiffHOas well as our baselines.
+B.4. Annotation Examples
+We provide a few examples of our textual annotations in Table 10. Please see our code release for the entire set of annotations***The code and data will be available here: https://diffh2o.github.io.
+B.5. Visualization Pipeline.
+We use AITViewer (Kaufmann et al., 2022) to visualize motions synthesized by DiffH2O as well as our baselines.
 IMoS synthesizes 15 frames of human-object interactions.
 They determine the hand which is in contact with the object according to GT, which can lead to a jump of the object from one hand to the other for certain actions.
-To make the frame rates comparable with DiffHO, we upsample the synthesis results of IMoS to 75 frames using spherical interpolation.
+To make the frame rates comparable with DiffH2O, we upsample the synthesis results of IMoS to 75 frames using spherical interpolation.
 Therefore, this jump may result in an artifact in which the object is floating from one hand to another.
@@
 We show however in this supplemental material (Table 7) that our model is still able to outperform IMoS even when all the GRAB data, including test data, is included in training the classifier.
-This is a setup which is more favorable for IMoS as the first frame of synthesized IMoS motion comes from test data.
+This is a setup which is more favorable for IMoS as the first frames of synthesized IMoS motion is taken from the test data.
 This re-confirms the ability of our method to encode diverse motion features that represents realistic hand-object interactions.
-| Method |
-DIV () | MM () | KID () | FID () | |||||
-Real Mocap | 1.0 | 1.0 | 1.13580.0164 | 0.31050.0163 | - | - | ||||
-| IMoS (Ghosh et al., 2023) | 0.7017 | 0.6754 | 1.10530.0128 | 0.28820.0114 | 0.005811 | 0.6267 | |||||
-| DiffHO | 0.8125 | 0.9028 | 1.14400.0115 | 0.31760.0127 | 0.005346 | 0.8342 | |||||
-Real Mocap | 1.0 | 1.0 | 1.13240.0119 | 0.23550.0044 | - | - | ||||
-| IMoS* (Ghosh et al., 2023) | 0.7161 | 0.8000 | 1.11730.0112 | 0.25130.0104 | 0.006038 | 0.7101 | |||||
-| DiffHO | 0.7836 | 0.8125 | 1.12250.0120 | 0.21740.0074 | 0.005494 | 0.8821 |
+Furthermore, we train two models, one that uses only hand 3D joints as input and another incorporating object 3D position as additional input.
+We demonstrate that DiffH2O’s synthesized object motions outperform IMoS in both settings.
+This confirms the benefits of jointly modeling hand-object motion instead of decoupling it into separate components.
+Note also that as expected, the classifier can correctly identify all test sequences of the real mocap data, since it has seen the test data during training in the IMoS setting.
 FID and KID metrics.
@@
 Table 8 also includes the standard deviations for the multimodality and diversity metrics based upon 20 repetitions of the evaluation.
-| Method | DIV () | MM () | KID () | FID () | ||||
-Ground-truth | 1.06320.0131 | 0.21550.0091 | - | - | |||
-| IMoS (Ghosh et al., 2023) | 1.02370.0158 | 0.25380.0103 | 0.008065 | 0.6267 | ||||
-| DiffHO | 1.07570.0129 | 0.30370.0079 | 0.006697 | 0.8342 | ||||
-Ground-truth | 1.07570.0129 | 0.20020.0063 | - | - | |||
-| IMoS* (Ghosh et al., 2023) | 1.04710.0129 | 0.22020.0093 | 0.008641 | 0.6593 | ||||
-| DiffHO | 1.09420.0125 | 0.23070.0060 | 0.007503 | 0.8425 |
+| Method |
+DIV () | MM () | KID () | FID () | |||
+| Unseen subject split | Real Mocap | 1.0 | 1.0 | 1.13580.0164 | 0.31050.0163 | - | - | ||
+| IMoS (Ghosh et al., 2023) | 0.7017 | 0.6754 | 1.10530.0128 | 0.28820.0114 | 0.005811 | 0.6267 | |||
+| DiffH2O | 0.8125 | 0.9028 | 1.14400.0115 | 0.31760.0127 | 0.005346 | 0.8342 | |||
+| Unseen object split | Real Mocap | 1.0 | 1.0 | 1.13240.0119 | 0.23550.0044 | - | - | ||
+| IMoS* (Ghosh et al., 2023) | 0.7161 | 0.8000 | 1.11730.0112 | 0.25130.0104 | 0.006038 | 0.7101 | |||
+| DiffH2O | 0.7836 | 0.8125 | 1.12250.0120 | 0.21740.0074 | 0.005494 | 0.8821 |
+| Method | DIV () | MM () | KID () | FID () | |
+| Unseen subject split | Real Mocap | 1.06320.0131 | 0.21550.0091 | - | - |
+| IMoS (Ghosh et al., 2023) | 1.02370.0158 | 0.25380.0103 | 0.008065 | 0.6267 | |
+| DiffH2O | 1.07570.0129 | 0.30370.0079 | 0.006697 | 0.8342 | |
+| Unseen object split | Real Mocap | 1.07570.0129 | 0.20020.0063 | - | - |
+| IMoS* (Ghosh et al., 2023) | 1.04710.0129 | 0.22020.0093 | 0.008641 | 0.6593 | |
+| DiffH2O | 1.09420.0125 | 0.23070.0060 | 0.007503 | 0.8425 |
 Full-sequence training.
-Given an initial pose of the grasp moment, IMoS only synthesizes post-grasp sequences.
-To compare our results to IMoS in Table 1 of the main paper, we evaluated our models only on the post-grasp sequence part of the synthesis.
-In Table 9, we also provide diversity, multimodality, FID, KID and action recognition scores computed over the full sequence including pre- and post-grasp stages.
-We compare our results against real motions from ground-truth and demonstrate that we can achieve scores similar to ground-truth, which demonstrates our capability to synthesize realistic hand-object interactions.
+Given an initial pose of the grasp moment, IMoS only synthesizes interaction sequences.
+To compare our results to IMoS in Table 1 of the main paper, we evaluated our models only on the interaction sequence part of the synthesis.
+In Table 9, we also provide diversity, multimodality, FID, KID and action recognition scores computed over the full sequence including grasp and interaction stages.
+We compare our results against motions from motion capture and observe that we can achieve scores similar to real mocap, which demonstrates our capability to synthesize realistic hand-object interactions.
 | Method |
@@
 | Real Mocap (Full-sequence) | 1.0 | 1.0 | 1.09830.0059 | 0.22350.0042 | - | - | ||||
-| DiffHO (Full-sequence) | 0.7720 | 0.8031 | 1.09480.0120 | 0.15550.0057 | 0.0067 | 0.7993 |
-C.2. Text Input Evaluation Experiments
-In Table 10, we provide further details for the text input to the diffusion model.
-We first compare IMoS and our model trained on the subject split and tested on the object split.
-Notably, our model outperforms IMOS by 9.2% in action correctness, a significant improvement.
-When trained and tested on our unseen object splits, we find that our model remains competitive with IMoS.
-However, we notice that our model struggles with classes like apple and train, where new motions are introduced that differ from those in the training set.
-Our model must learn to generalize from eating bananas to apples, and while it achieves this, its performance decreases to 81.8%.
-Similarly, for the train class, our model’s accuracy drops to 50% since the unique motions of playing with a train were not observed during training.
-Nevertheless, our model demonstrates the ability to generalize to different object sizes, as it generates motions corresponding to recognizable actions for varied sizes of cylinders, pyramids, and torus in the training set.
-Finally, we train and test our model with the augmented sentences we proposed for GRAB, leading to improved generalization, especially for the classes “train” and “apple”.
-These new sentences provide detailed motion descriptions, such as “The person picks up the train using their right hand, then rides it through the air close to the table surface, and finally puts it on the table using their right hand.” The augmented sentences enhance generalization and provide control over specifics like handedness, motion type, and physical location.
-Testing a model trained solely with baseline sentences on augmented sentences reveals a decrease in generalization, as anticipated.
-We observed a significant 38.7% performance gap between baseline and augmented sentences, showing the importance of text style.
-This highlights the limitations of relying solely on simple sentences for training and underscores the value of incorporating diverse and detailed descriptions to enhance model performance and generalization capabilities.
-C.3. Detailed Investigation User Study
-In our user study, we provide each user with 30 side-by-side comparisons.
-In our investigation, our model showed low realism but high diversity for the “pass” action, influenced by users perceiving “pass” action label in the GRAB dataset as transferring an object between hands, not to another person.
-In the “offhand” category, our model excelled in realism, while IMoS exhibited artifacts.
-This is due to the fact that IMoS determines the hand which is in contact with the object according to ground-truth, which leads to a jump of the object for the “off-hand” category.
-C.4. More Qualitative Results and Failure Cases
-We provide more qualitative results of our model with simple instructions in Fig.
-5(a), for grasp guidance in Fig.
-5(b), and with detailed instructions in Fig.
-5(c).
-Furthermore, we present three failure cases: the generated action does not match the text prompt (Fig.
-6(a)), the grasp reference is ignored during guidance (Fig.
-6(b)), or the hand in the text prompt is ignored (Fig.
-6(c)).
-| Split | Textual input form |
-Action correctness over object classes | ||||||||||||||||||
-| Method | Train split | Test split | Train input | Test input | Overall | apple | mug | train |
-elephant |
-| IMOS* | subject | object | simple | simple | 90.3 | 90.9 | 89.5 | 77.8 | 100 | 80.0 | 100 | 83.3 | 100 | ||||||||
-| Ours | subject | object | simple | simple | 99.5 | 100 | 97.4 | 100 | 100 | 100 | 100 | 100 | 100 | ||||||||
-| Ours | object | object | simple | simple | 89.7 | 81.8 | 92.1 | 50.0 | 100 | 100 | 100 | 100 | 100 | ||||||||
-| Ours | object | object | simple | comprehensive | 54.8 | 40.9 | 55.3 | 16.7 | 52.6 | 80.0 | 75 | 50.0 | 68.8 | ||||||||
-| Ours | object | object | comprehensive | comprehensive | 93.5 | 95.5 | 94.7 | 77.8 | 100 | 95.0 | 93.6 | 100 | 93.6 | ||||||||
+| DiffH2O (Full-sequence) | 0.7720 | 0.8031 | 1.09480.0120 | 0.15550.0057 | 0.0067 | 0.7993 |
+| Object Name | Annotation |
+| camera | The person uses their left hand to pick up the camera, brings it closer to their eyes to take a photo, and then places it back on the table with their right hand. |
+| stapler | The person repeatedly picks up and places the stapler on the table always with their right hand. |
+| mouse | The person picks up the mouse with their right hand, passes it to someone on their right below chest level, and then places it back on the table with their right hand. |
+| alarm clock | The person repeatedly picks up the alarm clock with their right hand, lifts it, and places it back on the table with their right hand. |
+| duck | Using the right hand, the person picks up the duck, examines it closely, and then places the duck back down on the table. |
+| bowl | The person picks up the bowl with both hands, drinks from it, and places it back on the table with both hands. |
+| headphones | The person picks up the headphones with their left hand, puts them on with both hands, and places them on the table with both hands. |
+| banana | The person picks up the banana with their right hand, passes it to their left hand, then hands it to someone on their left side at shoulder level with their left hand, and finally places it back on the table with their right hand.
+|
+C.2. Analysis on Generalizability
+When we analyze per-unseen-object performance on GRAB, our method effectively generalizes to unseen objects with various sizes and shapes (e.g., apple, clock, train, elephant, mug, etc.).
+We also evaluated the influence of object scale on three test objects by using objects of the same shape but different scales during training: our model performs well on medium cylinder and large torus (584cm³), but faces challenges on small pyramid (64cm³, video: 04.40, low contact-ratio).
+The average bounding-box volumes for our train and test sets are 741cm³ and 625cm³, respectively.
+In conclusion, our model adapts well to unseen shapes and a wide range of sizes, but largely out-of-distribution sizes (e.g., also large_cracker_box from HO3D: 2208cm³) lead to less favorable physics metrics.
+C.3. Discussion on Grasp References
+Here we provide a more detailed discussion about the role of grasp references that are used for guidance.
+Strengths:
+Grasp guidance is a technique to increase the controllability and improve generalization on unseen objects if grasp references are available.
+Note, however, that our model without guidance already significantly outperforms baselines (cf.
+Table 1 in the main paper), making it an ideal tool for potential downstream use cases.
+Grasp guidance yields better generalization (e.g., less interpenetration as shown in Table 2 in the main paper), because it provides a signal to the diffusion model about how an object should be grasped.
+Especially for unseen object geometry, this can serve as important information about feasible grasping poses of the hand.
+One of the other main advantages of grasp guidance is that it allows defining the grasping hand(s) on the object, which may be important if rich textual annotations such as the ones presented in Sec.
+4.4 are not available.
+Weaknesses and Failures:
+Grasp guidance faces an important trade-off: while the signal can be helpful for object generalization and controllability of handedness, it requires access to single hand-object poses.
+We detail in Section B.1 how such grasp references can be obtained.
+Hence, there are conversion steps that are necessary to make guidance work on unseen object geometry, which require some manual effort.
+For instance, due to the small dataset size of GRAB and the similar approaching direction for all objects, an alignment step is necessary to make the guidance signal useful.
+In the future, larger dataset distributions or additional data augmentation techniques may help alleviate this.
+Furthermore, as our model is conditioned on both the text description and the grasp guidance, one needs to make sure that the two align: for example, if the grasp reference shows grasping a cup from above, where the prompted action is ”drink from the cup”, the model is forced to select which signal to follow.
+In practice, as the guidance is only a weak guidance signal at a single frame, the model will favor the text description over the guidance.
+Another limiting factor is the size of the unseen objects as discussed in Section C.2.
+For example, if the object is too large, the target grasp may be reached in the grasping phase, but after imputing the grasping into the interaction stage, the model may neglect the out-of-distribution grasp and output a grasp that displays physical artifacts such as interpenetration.
+Overall, grasp guidance is a technique that shows the potential to provide a link between how an object should be grasped and subsequent actions that can be prompted via text.
+While there are limiting factors in its current state, i.e., manual alignment and matching text prompt with the grasp reference, future work could focus on enhancing existing datasets and using improved guidance techniques (Karunratanakul et al., 2024) to further increase performance.
 Appendix D Computational Resources
@@
 The total inference time for 32 samples is approximately 300 seconds.
+Appendix E Ethical Discussion
+Our method uses a generative model, namely diffusion models, to synthesize realistic hand-object motions.
+While the generated outputs at this point are in 3D and not rendered on photorealistic images with natural backgrounds, such a feature could be added in the future.
+This may lead to malicious use of our research, such as generating virtual deepfakes.
+Other downstream applications that could have problematic use cases are robotics, where our generated data could be used to train robots in handling objects in a desired way.
+For example, robots could be taught how to handle weapons.
+Moving forward with the development of our method, we hope that openly discussing the technical details, implementation, and sharing our code will ensure that the technology is well comprehended, accessible to the public and measures to prevent and flag malicious uses are easier to implement.